### v1

# Realistic Extreme Image Rescaling via Generative Latent Space Learning

Abstract
Image rescaling aims to learn the optimal downscaled low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart. This process is crucial for efficient image processing and storage, especially in the era of ultra-high definition media. However, extreme downscaling factors pose significant challenges due to the highly ill-posed nature of the inverse upscaling process, causing existing methods to struggle in generating semantically plausible structures and perceptually rich textures. In this work, we propose a novel framework called Latent Space Based Image Rescaling (LSBIR) for extreme image rescaling tasks. LSBIR effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model to generate realistic HR images. The rescaling is performed in the latent space of a pre-trained image encoder and decoder, which offers better perceptual reconstruction quality due to its stronger sparsity and richer semantics. LSBIR adopts a two-stage training strategy. In the first stage, a pseudo-invertible encoder-decoder models the bidirectional mapping between the latent features of the HR image and the target-sized LR image. In the second stage, the reconstructed features from the first stage are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details. Extensive experiments demonstrate the superiority of LSBIR over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/wwangcece/LSBIR.
1 Introduction
With the explosive growth of ultra high-resolution digital visual data, image rescaling has become essential for enhancing data storage and transmission efficiency, optimizing resource usage in real-time applications, and ensuring proper display on devices with varying resolutions. During image rescaling, high-resolution (HR) images are often downscaled to visually pleasing low-resolution (LR) images for transmission or display. Subsequently, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike the super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
Figure 1: Visual comparison of 16$\times$ image rescaling results. HCFlow struggles to obtain sufficient texture, while the reconstruction results of VQIR lack accurate semantic and structural information.
In recent years, various image rescaling methods have been proposed, which can be classified into two main architectures: encoder-decoder [11, 22] and invertible neural networks [30, 16]. The encoder-decoder architecture employs two sets of neural networks to perform downscaling and upscaling operations. In contrast, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling operations. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
However, with the advancement of HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR visual data. When dealing with extreme rescaling factors, existing methods often struggle to generate sufficient details. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [36] and BDFlow [13] utilize pre-trained StyleGAN [8, 9] to generate high-quality upscaling face images. VQIR [29], on the other hand, leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from shortcomings in overall semantics and structure. As shown in Fig. 1, while VQIR generates richer details compared to HCFlow, its reconstruction results lack precise semantic and structural information. This is due to the use of both latent space and pixel space losses. The introduction of pixel-space loss can force the model to compromise on both optimization paths, leading to the generation of significant artifacts.
Figure 2: Image rescaling within pixel-space and latent-space. Different from methods that impose pixel loss, our approach applies loss only in the latent space, thereby ensuring the semantic integrity of the reconstruction results.
Therefore, we propose a novel framework named Latent Space Based Image Rescaling (LSBIR) for extreme rescaling, which does not impose any pixel loss on the reconstruction results, as illustrated in Fig. 2. Rescaling operations in the latent space have a broader receptive field, allowing richer information to be embedded into the LR images. Additionally, training in the latent space preserves the semantic integrity of the resampled features [7], better aligning with the intrinsic priors of pre-trained generative models. Specifically, LSBIR adopts a two-stage training strategy. In the first stage, a simple convolutional neural network (CNN) is used to rescale the latent code of the HR images. Concurrently, an invertible neural network (INN) completes the bidirectional mapping between the feature domain and the pixel domain. In the second stage, the rescaled features from the first stage guide the generation process of the pre-trained Stable Diffusion (SD) model, enhancing the details of the rescaled features. The first stage ensures that the rescaled features retain complete semantic information of the HR images, while the second stage generates richer textures and improves perceptual quality, balancing fidelity and realism in the final results.
Our contributions can be summarized as follows:
  • We propose a Latent Space Based Image Rescaling framework for extreme rescaling. Without imposing any pixel loss on the reconstruction results, our method exhibits excellent performance in global semantic and structural reconstruction.
  • We propose a novel decoupled strategy for feature rescaling. By decoupling the feature-pixel transform operation from the rescaling operation, this module can enhance the fidelity of feature reconstruction.
  • To the best of our knowledge, we are the first to apply diffusion models to the field of extreme image rescaling. We demonstrate that the rich knowledge contained in the pre-trained generative diffusion model can significantly enhance the perceptual quality of rescaling results.
Figure 3: Overview of proposed Latent-Space Based Image Rescaling (LSBIR). In the first training stage, we train a feature rescaling network in the latent space of a pre-trained VAE, which outputs LR images and rescaled features. In the second training stage, we use the features from the first stage as conditions and train an adapter to guide the generation process of the pre-trained SD model.
2 Related work
2.1 Image Rescaling
Image rescaling aims to downscale HR images to visually coherent LR images and then plausibly reconstruct the original HR images. Unlike image SR, image rescaling involves joint modeling of both downscaling and upscaling processes, leading to higher reconstruction accuracy. Earlier works, such as those by [11, 15], utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes. Instead of directly generating LR images, CAR [22] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, thus avoiding pixel-level constraints on the downscaled images. To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task in recent years [14, 37, 33, 31]. IRN [30] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency components into a case-agnostic latent distribution for efficient reconstruction. HCFlow [16] assumes that high-frequency (HF) information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation. In the realm of extreme image rescaling, generative priors are usually employed to constrain the outcomes. Notably, GRAIN [36] and VQIR [29] leverage pre-trained StyleGAN [8, 9] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality. Nonetheless, these approaches are often limited in scope, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
Recently, the emergence of the pre-trained SD model [20] has showcased its efficacy in both content generation [32] and image restoration [25]. Despite this progress, to date, no effort has been made to integrate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model. The former is represented by SR3 [21], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [25], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios.
However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling. In this work, we decompose extremec image rescaling into two sub-tasks: feature rescaling in the latent space and perception-oriented optimization. By doing so, we could explore the potential application of diffusion models in this field.
3 Methodology
3.1 Overview of LSBIR
Our LSBIR architecture, shown in Fig. 3, involves two training stages: Feature Rescaling in the Latent Space and Perception-oriented Optimization. Initially, the input HR image is encoded into a latent feature $z$, which is then rescaled by the Decoupled Feature Rescaling Module (DFRM) to generate an LR image $y$ and a rescaled latent feature $\hat{z}$. The first stage aims to align $\hat{z}$ with $z$ while ensuring $y$ retains the content of the original HR image, preserving semantic information for the second stageś refinement. Although decoding $\hat{z}$ produces high-fidelity HR images, they lack texture detail. Thus, in the second stage, we use an adapter to guide the pre-trained SD model, generating more realistic textures. During inference, we could balance fidelity and realism by weighting $\hat{z}$ from both stages.
3.2 Feature Rescaling in the Latent Space
3.2.1 Overall Network Architecture
In the first stage, we perform rescaling of the HR features $z$ in the latent space, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$. Here, $G_{e}$ is an encoder used to downscale the input features to the target size, while $G_{d}$ represents the decoder used to reconstruct the original high-resolution features. Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{p}$ (1)
$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{p}$ (2)
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR [28] problem. Conversely, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [18]. Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ might negatively impact subsequent upscaling.
Considering the above issues, we devise the DFRM as shown in Fig. 4. Inspired by FGRN [14], we use a set of CNNs to perform downscaling and upscaling, and a separate set of INNs to perform the bidirectional mapping between the feature domain and the pixel domain. To achieve balance between feature rescaling accuracy and visual quality of LR images, we design another pixel guidance module in the encoder to incorporate pixel-level information. Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow w\rightarrow\hat{z}\right)$; $\left(w\leftrightarrow y\right)$ as follows:
$w=G_{e}\left(x,z\right)$ (3)
$w\approx F^{-1}\left(F\left(w\right)\right)=F^{-1}\left(y\right)$ (4)
$\hat{z}=G_{d}\left(w\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and reverse processes of the INN, respectively. Given the latent variable $z\in R^{H/8\times W/8\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $w\in R^{H/N\times W/N\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $w$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{H/N\times W/N\times 3}$ through quantization. Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $w$. Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
Figure 4: Detailed architecture of Decoupled Feature Rescaling Module (DFRM).
3.2.2 Domain Converter
The domain converter module consists of a series of stacked invertible blocks [3], which can be defined as $F=f_{1}\circ f_{2}\circ...f_{K}$. For the $i$-th invertible block $f_{i}$ in the forward process, the input $w_{k}$ is split into $w_{k}^{1}$ and $w_{k}^{2}$ along the channel axis, and they undergo affine coupling layers for information fusion as follows:
$w_{k+1}^{1}=w_{k}^{1}+\phi\left(w_{k}^{2}\right)$ (6)
$w_{k+1}^{2}=w_{k}^{2}\otimes\exp\left(\rho\left(w_{k+1}^{1}\right)\right)+\eta\left(w_{k+1}^{1}\right)$ (7)
Here, $\phi\left(\cdot\right)$, $\rho\left(\cdot\right)$ and $\eta\left(\cdot\right)$ could be arbitrary functions and we adopt the widely used RRDB [26] for its good representation ability. Finally, the output of the last invertible module, i.e., $f_{K}$, is quantized to obtain the LR image $y$. Similarly, in the reverse process, the LR image $y$ will be used as an input to reconstruct $w$. For the $i$-th invertible block $f_{i}$, the input $w_{k+1}$ is split into $w_{k+1}^{1}$ and $w_{k+1}^{2}$ and undergo affine coupling layers as follows, thus obtaining the low-resolution representation $w$:
$w_{k}^{2}=\left(w_{k+1}^{2}-\eta\left(w_{k+1}^{1}\right)\right)\otimes\exp\left(-\rho\left(w_{k+1}^{1}\right)\right)$ (8)
$w_{k}^{1}=w_{k+1}^{1}-\phi\left(w_{k}^{2}\right)$ (9)
3.2.3 Loss Function
First stage training involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $w$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow w\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow w\rightarrow y\rightarrow\hat{w}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{p}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{p}\end{split}$ (10)
This combination can facilitate the network to reconstruct $\hat{z}$ with high-fidelity under both the existence and absence of quantization operations. Additionally, considering that the guidance loss should be decoupled from the reconstruction loss as much as possible, we design it as follows:
$\mathcal{L}_{\mathrm{gui.}}=\parallel\mathrm{sg}\left(F\left(G_{e}\left(x,z\right)\right)\right)-\mathrm{Bicubic}(x)\parallel_{p}\$ (11)
Here, $\mathrm{sg}\left(\cdot\right)$ is the stop-gradient operation preventing the gradient flowing into $G_{e}$ or $G_{d}$, which decouples the feature-pixel domain bidirectional mapping and feature rescaling operations. Finally, the loss in the first stage is a combination of these two losses:
$\mathcal{L}_{\mathrm{stage1}}=\lambda_{1}\mathcal{L}_{\mathrm{rec.}}+\lambda_{2}\mathcal{L}_{\mathrm{gui.}}\$ (12)
Table 1: Quantitative comparisons with different methods at 16$\times$ and 32$\times$. The symbols $\uparrow$ and $\downarrow$ respectively represent that higher or lower values indicate better performance. Bold represents the best and underline represents the second best.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$
DIV2K EDSR 24.64 22.13 0.6572 0.6056 0.5376 0.6219 0.3162 0.4184 34.15 23.28 0.3397 0.4193
RCAN 24.34 22.06 0.6479 0.6034 0.5485 0.6307 0.3269 0.4349 31.57 23.18 0.3408 0.4091
ESRGAN 23.15 21.15 0.5946 0.5808 0.4478 0.5788 0.2378 0.4237 59.80 43.40 0.6161 0.3725
IRN 26.04 22.84 0.7020 0.6212 0.5153 0.6095 0.3041 0.4036 43.72 28.70 0.2799 0.2336
HCFlow 26.66 23.89 0.7176 0.6467 0.4885 0.5816 0.2866 0.3852 46.43 37.25 0.2735 0.2792
VQIR 23.91 22.02 0.6498 0.5823 0.3174 0.4568 0.1024 0.2663 64.04 58.21 0.6350 0.6293
Ours 22.58 20.41 0.6171 0.5442 0.3559 0.4763 0.1188 0.1964 67.73 65.27 0.6934 0.6528
Urban100 EDSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
RCAN 19.96 18.14 0.5017 0.4452 0.5859 0.6880 0.3745 0.4808 40.31 24.46 0.3561 0.4130
ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
IRN 22.15 18.93 0.6155 0.4772 0.5114 0.6567 0.3338 0.4457 58.73 32.03 0.3217 0.2210
HCFlow 22.59 19.86 0.6335 0.5185 0.4841 0.6048 0.3125 0.4101 60.62 47.94 0.3182 0.2827
VQIR 20.27 18.44 0.5782 0.4427 0.3038 0.5045 0.1418 0.3290 71.56 61.64 0.6619 0.6568
Ours 22.16 17.76 0.6809 0.4342 0.2647 0.5164 0.1253 0.2904 72.71 69.72 0.6832 0.6602
CLIC2020 EDSR 26.33 23.92 0.7408 0.7062 0.5109 0.5742 0.3057 0.4024 33.17 21.77 0.3392 0.4307
RCAN 26.37 24.04 0.7407 0.7072 0.5133 0.5777 0.3126 0.4151 31.59 22.07 0.3469 0.4180
ESRGAN 25.43 23.48 0.7018 0.6991 0.4213 0.5347 0.2209 0.4014 57.80 40.98 0.6107 0.3537
IRN 28.55 24.93 0.7907 0.7228 0.4676 0.5528 0.2746 0.3776 47.66 27.81 0.2950 0.2495
HCFlow 28.82 26.21 0.7949 0.7451 0.4542 0.5325 0.2635 0.3611 47.72 37.15 0.2733 0.2884
VQIR 25.69 24.12 0.7363 0.6942 0.2854 0.4191 0.0901 0.2445 64.36 54.83 0.6105 0.6068
Ours 24.62 22.59 0.7063 0.6551 0.3423 0.4618 0.1217 0.2036 66.48 61.67 0.6662 0.6229
DIV8K EDSR 26.56 24.48 0.6776 0.6472 0.5437 0.6048 0.3358 0.4325 30.83 22.30 0.3701 0.4686
RCAN 26.63 24.57 0.6778 0.6481 0.5470 0.6085 0.3430 0.4449 29.93 22.52 0.3796 0.4502
ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
IRN 28.95 25.46 0.7402 0.6636 0.4932 0.5870 0.3044 0.4066 42.61 27.76 0.3216 0.2741
HCFlow 29.22 26.69 0.7470 0.6857 0.4793 0.5616 0.2932 0.3906 42.71 35.53 0.2948 0.3128
VQIR 25.79 24.26 0.6889 0.6350 0.3221 0.4457 0.1066 0.2628 56.76 51.44 0.5769 0.6203
Ours 24.39 22.94 0.6515 0.5943 0.3846 0.5058 0.1524 0.2451 58.28 55.68 0.6663 0.6231
3.3 Perception-oriented Optimization
The $\hat{z}$ in the first stage can be passed to the decoder $\mathcal{D}$ to obtain the rescaled HR image. Although this result has high fidelity, it lacks rich texture. Therefore, training in the second stage is designed to enhance perceptual quality. Inspired by relevant content generation [32, 19], we adopt the adapter architecture to guide the generation of the SD model. The adapter module takes the rescaled features $\hat{z}$ as input and extracts multi-scale features in a coarse-to-fine manner, allowing it to output multi-scale features of different sizes:
$\{F_{c}^{1},F_{c}^{2},...,F_{c}^{K}\}=\mathrm{Ada}(\hat{z})$ (13)
Subsequently, these features are element-wisely added to different layers of the denoising U-Net. Considering that image rescaling task places a higher emphasis on fidelity, we employ conditional control over both the encoding and decoding layers of the denoising U-Net:
$F_{\mathrm{enc}}^{i}=F_{\mathrm{enc}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (14)
$F_{\mathrm{dec}}^{i}=F_{\mathrm{dec}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (15)
Here, $F_{\mathrm{enc}}^{i}$ and $F_{\mathrm{dec}}^{i}$ represent the output features of the encoder or decoder of the denoising U-Net at the $i$-th layer. Finally, the SD model and adapter are optimized using the following loss function:
$\mathcal{L_{\mathrm{stage2}}}=\mathrm{\mathbb{E}}_{\left(z,\epsilon,\hat{z},t\right)}\parallel\epsilon-\epsilon_{\theta}\left(z_{t},t,\hat{z}\right)\parallel^{2}$ (16)
Figure 5: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results. Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
Figure 6: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both HR images and corresponding edge maps. Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics
We train the proposed model using the DIV2K dataset [1] and the Flickr2K dataset [17]. Due to the low resolution of conventional benchmark datasets such as Set5, Set14, and BSD100, they are not suitable for extreme rescaling tasks [5]. Therefore, we utilized four high-resolution datasets: DIV2K, Urban100 [6], CLIC2020 [23], and DIV8K [5] to evaluate our model. We adopt PSNR and SSIM [27] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [34] and DISTS [2] are used. For non-reference image quality assessment, we use metrics including MUSIQ [10] and CLIPIQA [24].
4.1.2 Training Details
Our LSDIR is built based on the SD 2.1-base model and the training is divided into two stages. First, we train the DFRM together with the pre-trained VAE model for 200K iterations using Eq. 12 The learning rate is initialized as $2\times 10^{-4}$ and reduced by half every 50k iterations. After the training of DFRM, we adopt Eq. (16) to train the Adapter and finetune the diffusion U-Net. We only finetune the linear layers of the U-Net for its better performance. During both stages, we adopt Adam optimizer [12] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and set the mini-batch size to 32.
4.2 Results and Comparison with SOTA
[…]
4.2.1 Quantitative Results
Tab. 1 presents the 16$\times$ and 32$\times$ quantitative image rescaling results on four datasets. The performance of SR methods are generally poor. While methods based on INNs achieve the highest PSNR and SSIM scores due to consideration of the correlation between downscaling and upscaling. VQIR exhibits the best LPIPS by fully utilizing the prior provided by pre-trained VQGAN and being trained based on LPIPS-loss. As our proposed method is trained in the latent space, the results exhibit the best semantic integrity, which leads to significant improvements in two non-reference metrics. It is important to note that while our method slightly lags behind VQIR in terms of certain full-reference metrics, it excels in reconstructing accurate semantics and regular structures, as will be demonstrated in the following section.
4.2.2 Qualitative Results
In Fig. 5, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors. Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as EDSR, RCAN, IRN, and HCFlow produce overly-smoothed results. Although ESRGAN incorporates perceptual optimization it still cannot generate adequate details. VQIR leverages the pre-trained VQGAN as a prior, allowing it to generate rich details and sharp edges. However, it struggles with faces and more complex structures. Thanks to the powerful prior provided by the pre-trained SD and the rich semantic information contained in the latent space of the pre-trained VAE, our method effectively addresses these issues. It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as faces, text and architectural structures.
The quantitative experiments demonstrate that our model exhibits outstanding performance on the Urban100 dataset, confirming that our method produces more accurate structural features. Therefore, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 6. Traditional regression models, such as EDSR, RCAN, IRN, and HCFlow, tend to generate overly-smooth results with few sharp edges. Although GAN-based methods, including ESRGAN and VQIR, generate more details, their edge information lacks accuracy and regularity. In contrast, our approach, by leveraging the prior knowledge encapsulated in the Stable Diffusion model, is able to reconstruct more accurate and rich edge details.
4.3 Ablation Study
We conduct ablation experiments to validate the effectiveness of each component. All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K.
4.3.1 Effectiveness of Feature Rescaling
To verify the effectiveness of the feature rescaling operation, we move the rescaling operation from the latent space to the pixel space. This means we directly train a downsampling network to downscale the HR $x$ to LR $y$. Subsequently, $y$ is used as a condition to guide the generation process of the SD model. As shown the 1st row in Tab. 2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by lower LPIPS and DISTS.
In Fig. 7, we plot the LPIPS variation curves on the validation set for the two rescaling schemes. The pixel space-based rescaling scheme requires a longer time to converge, whereas the latent space based scheme can stably converge to a better result in a shorter time. Additionally, we observe that the pixel space based scheme experiences severe color shifts in the middle of training, as illustrated in Fig. 8. It requires a substantial number of iterations to learn the global color distribution of the reconstructed images. This is because using pixels as the conditioning input for Stable Diffusion makes it difficult to align with the intrinsic priors of the pre-trained model, leading to issues such as unstable training and longer training time.
Figure 7: The validation LPIPS curve for rescaling models based on latent space and pixel space.
Figure 8: Reconstructed results after 60k training steps. The pixel space based scheme shows severe color shifts.
Table 2: Ablation study of rescaling space and SD prior.
Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
$\times$ $\checkmark$ $\checkmark$ 0.3818 0.1333
$\checkmark$ $\times$ $\times$ 0.4675 0.3109
$\checkmark$ $\times$ $\checkmark$ 0.3559 0.1188
4.3.2 Effectiveness of SD Prior
To validate that SD model provides rich priors about real-world HR images, we directly omit the second training stage. As shown in Tab. 2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results.
4.3.3 Effectiveness of DFRM Design
Our DFRM includes three specific designs: the domain converter module, the pixel guidance module, and the stop-gradient operation in the loss function Equ. 11. To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig. 10. It can be observed that the introduction of the INN and the stop-gradient operation can both reduce reconstruction error. Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as shown in Tab. 3. Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
Ideally, the domain converter composed of a set of invertible neural networks should be able to accurately perform bidirectional mapping between the feature domain and the pixel domain, enabling the LR to share the same content with HR images. However, due to the limited representation power of INN, the intermediate features $w$ obtained by simply downscaling the latent code $z$ are not well-suited for conversion to the pixel domain. The introduction of pixel guidance allows for embedding some pixel-domain information into the intermediate features $w$, thereby achieving a balance between reconstruction performance and LR image quality. As shown in Fig. 9, the LR images exhibit significant noise and color distortions in the absence of pixel guidance. However, the introduction of pixel guidance significantly alleviates these issues.
Figure 9: Comparisons of LR images visual quality with and without pixel guidance.
Table 3: Ablation study of pixel guidance module. The introduction of pixel guidance can significantly enhance the quality of LR images.
Method PSNR\ $\uparrow$ SSIM\ $\uparrow$ NIQE\ $\downarrow$
w/o pixel guidance 19.88 0.5898 22.33
with pixel guidance 21.99 0.7237 13.27
Figure 10: The validation loss curves when removing a specific component from the DFRM.
4.4 Discussions
4.4.1 Comparison with JPEG Compression
Our model can be combined with existing image compression methods to achieve more extreme compression ratios. To validate this, we use the lossless compression method PNG to compress LR images and compare its performance with JPEG compression at various LR sizes. Specifically, we train rescaling models at three different scales (8$\times$, 16$\times$, and 32$\times$) and record the bpp by the LR images. We also adjust the quality parameters of JPEG compression accordingly for comparison. As shown in Fig. 11, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of JPEG compression. In Fig. 12, we present the visual quality, the bpp and the LPIPS of the reconstructed images. Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics. Additionally, JPEG compression tends to produce noticeable block artifacts and color distortions in the background regions of the images, whereas our model maintains superior reconstruction quality across all three scales. These comparisons demonstrate the potential of our model for image compression.
Figure 11: R-D curves of JPEG compression and our proposed model on the DIV2K validation set.
Figure 12: Visual comparisons between our model and JPEG compression. Below the images, the bpp and LPIPS of the reconstructed images are provided.
4.4.2 Fidelity-realism Trade-off
Our two-stage model outputs two features in the latent space ($\hat{z}$ and $\hat{z}_{0}$), which represent reconstruction results with high fidelity and high realism respectively. It is important to note that the high fidelity mentioned here does not refer to higher pixel fidelity but to semantic fidelity in the latent space. By fusing the results of the two stages, we can achieve a trade-off between fidelity and realism:
$z_{\mathrm{fus.}}=\alpha\hat{z}+\left(1-\alpha\right)\hat{z}_{0}$ (17)
$x_{\mathrm{fus.}}=\mathcal{D}\left(z_{\mathrm{fus.}}\right)$ (18)
As shown in Fig. 13, by adjusting the value of $\alpha$, we can achieve a smooth transition between fidelity and realism. More results are provided in the supplementary materials. As shown in Fig. 14, It can be observed that by adjusting the fusion weights, we can achieve a smooth transition between fidelity and realism.
Figure 13: By adjusting the fusion ratio of two latent codes, we can achieve a trade-off between fidelity and realism.
Figure 14: By adjusting the fusion weights, our model can achieve a trade-off between fidelity and realism. The perceptually optimized results display richer details, but they also exhibit some artifacts.
Figure 15: Visual comparisons of LR images.
5 Limitations
Although our model demonstrates superior performance in extreme image rescaling tasks, it still has some limitations. One of them is the quality of the LR images. As shown in Fig. 15, the LR images generated by IRN closely resemble the bicubic downsampled image. In contrast, the LR images produced by HCFlow and VQIR exhibit some color shifts. LR images generated by our model are slightly noisy. In the future, more advanced invertible network architectures and more effective loss functions could be explored to further improve the quality of LR images.
6 Conclusion
Extreme image rescaling is a highly challenging yet significant task. In this paper, we propose a novel framework called LSBIR to address issues such as insufficient texture and semantic inaccuracies. Specifically, we leverage the rich priors encapsulated in the pre-trained SD model to mitigate the ill-posed nature of rescaling. In the first stage, we design a DFRM to achieve bidirectional mapping between HR features and LR images. In the second stage, the rescaled features from the first stage serve as conditions to guide the generation process of the SD model through an effective adapter. Both qualitative and quantitative experiments demonstrate the superiority of our approach.

### v2

@@
-Realistic Extreme Image Rescaling via Generative Latent Space Learning
+One-step Generative Diffusion for Realistic Extreme Image Rescaling
 Abstract
-Image rescaling aims to learn the optimal downscaled low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart.
-This process is crucial for efficient image processing and storage, especially in the era of ultra-high definition media.
-However, extreme downscaling factors pose significant challenges due to the highly ill-posed nature of the inverse upscaling process, causing existing methods to struggle in generating semantically plausible structures and perceptually rich textures.
-In this work, we propose a novel framework called Latent Space Based Image Rescaling (LSBIR) for extreme image rescaling tasks.
-LSBIR effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model to generate realistic HR images.
-The rescaling is performed in the latent space of a pre-trained image encoder and decoder, which offers better perceptual reconstruction quality due to its stronger sparsity and richer semantics.
-LSBIR adopts a two-stage training strategy.
-In the first stage, a pseudo-invertible encoder-decoder models the bidirectional mapping between the latent features of the HR image and the target-sized LR image.
-In the second stage, the reconstructed features from the first stage are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details.
-Extensive experiments demonstrate the superiority of LSBIR over previous methods in both quantitative and qualitative evaluations.
-The code will be available at: https://github.com/wwangcece/LSBIR.
+Image rescaling aims to learn the optimal low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart, providing an efficient image processing and storage method for ultra-high definition media.
+However, extreme downscaling factors pose significant challenges to the upscaling process due to its highly ill-posed nature, causing existing image rescaling methods to struggle in generating semantically correct structures and perceptual friendly textures.
+In this work, we propose a novel framework called One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model.
+Specifically, OSIRDiff adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image.
+Then, the rescaled features are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details.
+The entire model is end-to-end trained to enable the diffusion priors to guide the rescaling process.
+Considering the spatially non-uniform reconstruction quality of the rescaled latent features, we propose a novel time-step alignment strategy, which can adaptively determine the generative strength of the diffusion model based on the degree of latent reconstruction errors.
+Extensive experiments demonstrate the superiority of OSIRDiff over previous methods in both quantitative and qualitative evaluations.
+The code will be available at: https://github.com/xxx/xxx.
 1 Introduction
-With the explosive growth of ultra high-resolution digital visual data, image rescaling has become essential for enhancing data storage and transmission efficiency, optimizing resource usage in real-time applications, and ensuring proper display on devices with varying resolutions.
-During image rescaling, high-resolution (HR) images are often downscaled to visually pleasing low-resolution (LR) images for transmission or display.
-Subsequently, an upscaling process is employed to restore the downscaled images to their original resolution.
-Unlike the super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
+With the explosive growth of ultra high-resolution (HR) images, image rescaling has become essential for enhancing data storage and transmission efficiency [44], optimizing resource usage in real-time applications [26], and ensuring proper display on devices with varying resolutions [25].
+During image rescaling, HR images are downscaled to low-resolution (LR) images for storage, transmission, or display.
+When HR images are needed, an upscaling process is employed to restore the downscaled images to their original resolution.
+Unlike image super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
 Figure 1: Visual comparison of 16$\times$ image rescaling results.
-HCFlow struggles to obtain sufficient texture, while the reconstruction results of VQIR lack accurate semantic and structural information.
-In recent years, various image rescaling methods have been proposed, which can be classified into two main architectures: encoder-decoder [11, 22] and invertible neural networks [30, 16].
-The encoder-decoder architecture employs two sets of neural networks to perform downscaling and upscaling operations.
-In contrast, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling operations.
+Recently published image rescaling methods can be classified into two main categories: encoder-decoder architecture [14, 31] and invertible neural networks [41, 18].
+The encoder-decoder architecture employs two neural networks to perform downscaling and upscaling.
+While, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling.
 Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
-However, with the advancement of HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR visual data.
-When dealing with extreme rescaling factors, existing methods often struggle to generate sufficient details.
+However, with the advancement of ultra HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR data [53, 8].
+Existing image rescaling methods often struggle to generate sufficient detail under extreme rescaling factors.
 Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results.
-GRAIN [36] and BDFlow [13] utilize pre-trained StyleGAN [8, 9] to generate high-quality upscaling face images.
-VQIR [29], on the other hand, leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images.
-Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from shortcomings in overall semantics and structure.
+GRAIN [53] and BDFlow [15] use pre-trained StyleGAN [10, 11] to generate high-quality upscaling face images.
+VQIR [39] leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images.
+Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from incorrect semantics and structure.
 As shown in Fig.
-1, while VQIR generates richer details compared to HCFlow, its reconstruction results lack precise semantic and structural information.
-This is due to the use of both latent space and pixel space losses.
-The introduction of pixel-space loss can force the model to compromise on both optimization paths, leading to the generation of significant artifacts.
-Figure 2: Image rescaling within pixel-space and latent-space.
-Different from methods that impose pixel loss, our approach applies loss only in the latent space, thereby ensuring the semantic integrity of the reconstruction results.
-Therefore, we propose a novel framework named Latent Space Based Image Rescaling (LSBIR) for extreme rescaling, which does not impose any pixel loss on the reconstruction results, as illustrated in Fig.
-2.
-Rescaling operations in the latent space have a broader receptive field, allowing richer information to be embedded into the LR images.
-Additionally, training in the latent space preserves the semantic integrity of the resampled features [7], better aligning with the intrinsic priors of pre-trained generative models.
-Specifically, LSBIR adopts a two-stage training strategy.
-In the first stage, a simple convolutional neural network (CNN) is used to rescale the latent code of the HR images.
-Concurrently, an invertible neural network (INN) completes the bidirectional mapping between the feature domain and the pixel domain.
-In the second stage, the rescaled features from the first stage guide the generation process of the pre-trained Stable Diffusion (SD) model, enhancing the details of the rescaled features.
-The first stage ensures that the rescaled features retain complete semantic information of the HR images, while the second stage generates richer textures and improves perceptual quality, balancing fidelity and realism in the final results.
+1, although VQIR generates richer details compared to HCFlow, it still fails to reconstruct accurate text and image patterns.
+To address the aforementioned issues, we propose a novel framework named One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images.
+Specifically, first, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image.
+Then, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality.
+Finally, the entire model is trained end-to-end to allow the diffusion priors to guide the rescaling process.
+Due to the spatially non-uniform reconstruction quality of the rescaled latent features, different regions of the image require varying levels of generative capability for restoration.
+Previous studies have shown that the time-step in diffusion models directly controls the generative capacity of the model [35].
+Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby achieving a balance between fidelity and realism.
+As shown in Fig.
+1, our model significantly enhances the overall quality of the rescaled images compared to other methods.
 Our contributions can be summarized as follows:
-  • We propose a Latent Space Based Image Rescaling framework for extreme rescaling.
-Without imposing any pixel loss on the reconstruction results, our method exhibits excellent performance in global semantic and structural reconstruction.
+  • We propose a One-Step Image Rescaling Diffusion framework that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
   • We propose a novel decoupled strategy for feature rescaling.
-By decoupling the feature-pixel transform operation from the rescaling operation, this module can enhance the fidelity of feature reconstruction.
-  • To the best of our knowledge, we are the first to apply diffusion models to the field of extreme image rescaling.
-We demonstrate that the rich knowledge contained in the pre-trained generative diffusion model can significantly enhance the perceptual quality of rescaling results.
-Figure 3: Overview of proposed Latent-Space Based Image Rescaling (LSBIR).
-In the first training stage, we train a feature rescaling network in the latent space of a pre-trained VAE, which outputs LR images and rescaled features.
-In the second training stage, we use the features from the first stage as conditions and train an adapter to guide the generation process of the pre-trained SD model.
+It decouples the generation of LR images from the rescaling operation, thereby improving the reconstruction accuracy of the rescaling process.
+  • We design a time-step alignment strategy to address the spatially non-uniform reconstruction quality of the rescaled latent features, allowing for more effective utilization of the prior information from the text-to-image diffusion model.
+Figure 2: Overview of the proposed One-Step Image Rescaling Diffusion (OSIRDiff).
+First, the HR images $x$ are encoded to obtain the latent features $z$.
+These latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$.
+Next, the rescaled latent features $\hat{z}$ are passed into a pre-trained diffusion model to perform a single denoising step for perceptual enhancement to obtain $\hat{z}_{0}$.
+Finally, the perceptually enhanced latent features $\hat{z}_{0}$ are decoded to obtain the rescaled image $\hat{x}$.
 2 Related Work
@@
 Unlike image SR, image rescaling involves joint modeling of both downscaling and upscaling processes, leading to higher reconstruction accuracy.
-Earlier works, such as those by [11, 15], utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes.
-Instead of directly generating LR images, CAR [22] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, thus avoiding pixel-level constraints on the downscaled images.
-To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task in recent years [14, 37, 33, 31].
-IRN [30] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency components into a case-agnostic latent distribution for efficient reconstruction.
-HCFlow [16] assumes that high-frequency (HF) information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation.
-In the realm of extreme image rescaling, generative priors are usually employed to constrain the outcomes.
-Notably, GRAIN [36] and VQIR [29] leverage pre-trained StyleGAN [8, 9] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality.
-Nonetheless, these approaches are often limited in scope, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
-Recently, the emergence of the pre-trained SD model [20] has showcased its efficacy in both content generation [32] and image restoration [25].
-Despite this progress, to date, no effort has been made to integrate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
+Earlier researches [14, 17] utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes.
+Instead of directly generating LR images, CAR [31] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, avoiding pixel-level constraints on the downscaled images.
+To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task [16, 54, 48, 43].
+IRN [41] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency (HF) components into a case-agnostic latent distribution for efficient reconstruction.
+HCFlow [18] assumes that the HF information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation.
+In extreme image rescaling, generative priors are usually employed to constrain the outcomes.
+GRAIN [53] and VQIR [39] leverage pre-trained StyleGAN [10, 11] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality.
+Nonetheless, these approaches are often limited in image contents, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
+Recently, the emergence of the pre-trained Stable Diffusion (SD) model [27] has showcased its efficacy in both content generation [47] and image restoration [34].
+Despite this progress, to date, no effort has been made to investigate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
 2.2 Diffusion Model for Image SR
@@
 However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling.
-In this work, we decompose extremec image rescaling into two sub-tasks: feature rescaling in the latent space and perception-oriented optimization.
-By doing so, we could explore the potential application of diffusion models in this field.
+In this work, we decompose extreme image rescaling into two sub-tasks: feature rescaling in the latent space and one-step perception enhancement with the pre-trained SD U-Net.
+2.3 Diffusion Acceleration
+To accelerate the inference process of diffusion models, researchers have proposed various optimization strategies.
+Fast diffusion samplers [30, 12, 21, 51] effectively reduce the sampling steps required by pre-trained diffusion models, lowering them from the initial thousand steps to as few as 20–50 steps.
+However, further reductions often lead to substantial performance degradation.
+Distillation techniques [23, 29, 42, 52, 45, 24] have shown significant promise in accelerating inference.
+For example, Progressive Distillation [29] minimizes sampling steps to under 10 by training a sequence of student models with progressively increasing stride.
+Instaflow [20] advances this further, achieving one-step generation using reflow and distillation methods.
+Recently, score-based distillation approaches, including Distribution Matching Distillation [45] and Variational Score Distillation [24], have utilized pre-trained text-to-image diffusion models as distribution matching losses to enable one-step text-to-image generation.
+Although these methods have made significant strides, high training costs remain a challenge.
+Inspired by Diffusion2GAN [9], we propose a low-overhead training approach based on Low-Rank Adaptation (LoRA) [6], achieving one-step acceleration for diffusion models and adaptability for image rescaling tasks.
 3 Methodology
-3.1 Overview of LSBIR
-Our LSBIR architecture, shown in Fig.
-3, involves two training stages: Feature Rescaling in the Latent Space and Perception-oriented Optimization.
-Initially, the input HR image is encoded into a latent feature $z$, which is then rescaled by the Decoupled Feature Rescaling Module (DFRM) to generate an LR image $y$ and a rescaled latent feature $\hat{z}$.
-The first stage aims to align $\hat{z}$ with $z$ while ensuring $y$ retains the content of the original HR image, preserving semantic information for the second stageś refinement.
-Although decoding $\hat{z}$ produces high-fidelity HR images, they lack texture detail.
-Thus, in the second stage, we use an adapter to guide the pre-trained SD model, generating more realistic textures.
-During inference, we could balance fidelity and realism by weighting $\hat{z}$ from both stages.
+3.1 Overview of OSIRDiff
+As illustrated in Fig.
+2, OSIRDiff consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, One-Step Perception Enhancement and Latent Decoding.
+First, the input HR images $x$ are mapped to the latent features $z$ via the pre-trained VAE encoder $\mathcal{E}$.
+Subsequently, the proposed Decoupled Feature Rescaling Module (DFRM) rescales $z$ to the target size, outputting LR images $y$ and the rescaled latent features $\hat{z}$.
+Our aim is to make $\hat{z}$ as close as possible to $z$ so that the rescaled results align with the latent space of the pre-trained VAE, while ensuring that $y$ and the original HR images $x$ share the same content.
+Then, the rescaled features $\hat{z}$ are fed into a denoising U-Net network to perform a single denoising step to get the perceptually enhanced latent features $\hat{z}_{0}$.
+Finally, perceptually enhanced latent features $\hat{z}_{0}$ are decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$.
+To dynamically adjust the generative capability based on the reconstruction quality of the rescaled latent features, we design a time-step alignment strategy, which consists of a time-step prediction module (TPM) and a refiner module.
 3.2 Feature Rescaling in the Latent Space
-3.2.1 Overall Network Architecture
-In the first stage, we perform rescaling of the HR features $z$ in the latent space, obtaining LR images $y$ and rescaled features $\hat{z}$.
+3.2.1 Network Architecture.
+In this part, we perform rescaling of the HR latent features $z$, obtaining LR images $y$ and rescaled features $\hat{z}$.
 A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$.
-Here, $G_{e}$ is an encoder used to downscale the input features to the target size, while $G_{d}$ represents the decoder used to reconstruct the original high-resolution features.
+Here, $G_{e}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{d}$ represents the decoder used to reconstruct the original HR features.
 Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
-$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{p}$ (1)
-$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{p}$ (2)
+$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{1}$ (1)
+$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{1}$ (2)
 However, these two losses are contradictory to each other.
-When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR [28] problem.
-Conversely, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [18].
-Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ might negatively impact subsequent upscaling.
+When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [38], and little information regarding the downscaling process can be embedded into the LR image.
+In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [22], which results in the LR image lacking any visual content.
+Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ could adversely impact subsequent upscaling.
+Figure 3: Detailed architecture of Decoupled Feature Rescaling Module (DFRM).
 Considering the above issues, we devise the DFRM as shown in Fig.
+3, which divides the rescaling operation into two sub-tasks: feature rescaling and LR image generation.
+Inspired by FGRN [16], we use a set of CNNs to perform downscaling and upscaling, and a separate INN to perform the bidirectional mapping between the feature domain and the pixel domain.
+To achieve a balance between feature rescaling accuracy and visual quality of LR images, we introduce another pixel guidance module in the encoder to incorporate pixel-level information, which can improve the quality of the LR image.
+Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}\right)$; $\left(z_{\text{lr}}\leftrightarrow y\right)$ as follows:
+$z_{\text{lr}}=G_{e}\left(x,z\right)$ (3)
+$z_{\text{lr}}\approx F^{-1}\left(F\left(z_{\text{lr}}\right)\right)=F^{-1}\left(y\right)$ (4)
+$\hat{z}=G_{d}\left(z_{\text{lr}}\right)$ (5)
+Here, $F$ and $F^{-1}$ represent the forward and backward processes of the INN, respectively.
+Given the latent variable $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $z_{\text{lr}}\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor.
+Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization.
+Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$.
+Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
+3.2.2 Loss Function.
+Training of this part involves two losses: reconstruction loss and guidance loss.
+Considering the information lost in quantization operations, the bidirectional mapping between $z_{\text{lr}}$ and $y$ is not exactly reversible.
+To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{y}_{lr}\rightarrow\hat{z}$, which can be expressed as the following formula:
+$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
+This combination can facilitate the network to reconstruct $\hat{z}$ with high fidelity in both the existence and absence of quantization operations.
+Additionally, the guidance loss is designed as follows:
+$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{e}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
+Finally, the loss function of this part is a combination of these two losses:
+$\mathcal{L}_{\mathrm{res.}}=\lambda_{\mathrm{rec.}}\mathcal{L}_{\mathrm{rec.}}+\lambda_{\mathrm{gui.}}\mathcal{L}_{\mathrm{gui.}}\$ (8)
+Figure 4: The rescaling operation leads to spatially non-uniform reconstruction quality of the rescaled latent features, which depends on the content of the image and the rescaling factor.
+3.3 One-Step Perception Enhancement
+The latent variable $\hat{z}$ obtained in Sec.
+3.2 retains sufficient semantic information about the HR image, but lacks adequate texture.
+Therefore, we are proposed to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part.
+Considering the original diffusion model requires a multi-step iterative sampling algorithm, this significantly increases inference time and makes it difficult to scale to practical applications.
+Inspired by recent work on accelerating diffusion models [9], we adopt an one-step inference strategy.
+Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ to perform a single denoising step as follows:
+$\hat{z}_{0}=\frac{\hat{z}-\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}$ (9)
+where $\hat{z}_{0}$ represents the perceptually enhanced latent features, $\bar{\alpha}_{t}$ is the noise scheduler of the forward diffusion process of the SD model, and $t$ represents the time-step associated with the noise intensity.
+Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}$:
+$\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$ (10)
+Then, we can adopt the loss function between $\hat{x}$ and $x$ to optimize the U-Net using LoRA layers.
+It should be noted that one of the key designs is the selection of the time-step $t$.
+Previous studies [35] have shown that the time-step in diffusion models could determine the trade-off between realism and fidelity.
+A larger $t$ results in stronger diversity and greater realism, but at the cost of fidelity.
+However, we have observed that the quality of the rescaled latent features exhibit non-uniform characteristics, as shown in Fig.
 4.
-Inspired by FGRN [14], we use a set of CNNs to perform downscaling and upscaling, and a separate set of INNs to perform the bidirectional mapping between the feature domain and the pixel domain.
-To achieve balance between feature rescaling accuracy and visual quality of LR images, we design another pixel guidance module in the encoder to incorporate pixel-level information.
-Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow w\rightarrow\hat{z}\right)$; $\left(w\leftrightarrow y\right)$ as follows:
-$w=G_{e}\left(x,z\right)$ (3)
-$w\approx F^{-1}\left(F\left(w\right)\right)=F^{-1}\left(y\right)$ (4)
-$\hat{z}=G_{d}\left(w\right)$ (5)
-Here, $F$ and $F^{-1}$ represent the forward and reverse processes of the INN, respectively.
-Given the latent variable $z\in R^{H/8\times W/8\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $w\in R^{H/N\times W/N\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor.
-Subsequently, a set of INNs perform the forward process to transform $w$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{H/N\times W/N\times 3}$ through quantization.
-Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $w$.
-Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
-Figure 4: Detailed architecture of Decoupled Feature Rescaling Module (DFRM).
-3.2.2 Domain Converter
-The domain converter module consists of a series of stacked invertible blocks [3], which can be defined as $F=f_{1}\circ f_{2}\circ...f_{K}$.
-For the $i$-th invertible block $f_{i}$ in the forward process, the input $w_{k}$ is split into $w_{k}^{1}$ and $w_{k}^{2}$ along the channel axis, and they undergo affine coupling layers for information fusion as follows:
-$w_{k+1}^{1}=w_{k}^{1}+\phi\left(w_{k}^{2}\right)$ (6)
-$w_{k+1}^{2}=w_{k}^{2}\otimes\exp\left(\rho\left(w_{k+1}^{1}\right)\right)+\eta\left(w_{k+1}^{1}\right)$ (7)
-Here, $\phi\left(\cdot\right)$, $\rho\left(\cdot\right)$ and $\eta\left(\cdot\right)$ could be arbitrary functions and we adopt the widely used RRDB [26] for its good representation ability.
-Finally, the output of the last invertible module, i.e., $f_{K}$, is quantized to obtain the LR image $y$.
-Similarly, in the reverse process, the LR image $y$ will be used as an input to reconstruct $w$.
-For the $i$-th invertible block $f_{i}$, the input $w_{k+1}$ is split into $w_{k+1}^{1}$ and $w_{k+1}^{2}$ and undergo affine coupling layers as follows, thus obtaining the low-resolution representation $w$:
-$w_{k}^{2}=\left(w_{k+1}^{2}-\eta\left(w_{k+1}^{1}\right)\right)\otimes\exp\left(-\rho\left(w_{k+1}^{1}\right)\right)$ (8)
-$w_{k}^{1}=w_{k+1}^{1}-\phi\left(w_{k}^{2}\right)$ (9)
-3.2.3 Loss Function
-First stage training involves two losses: reconstruction loss and guidance loss.
-Considering the information lost in quantization operations, the bidirectional mapping between $w$ and $y$ is not exactly reversible.
-To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow w\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow w\rightarrow y\rightarrow\hat{w}\rightarrow\hat{z}$, which can be expressed as the following formula:
-$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{p}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{p}\end{split}$ (10)
-This combination can facilitate the network to reconstruct $\hat{z}$ with high-fidelity under both the existence and absence of quantization operations.
-Additionally, considering that the guidance loss should be decoupled from the reconstruction loss as much as possible, we design it as follows:
-$\mathcal{L}_{\mathrm{gui.}}=\parallel\mathrm{sg}\left(F\left(G_{e}\left(x,z\right)\right)\right)-\mathrm{Bicubic}(x)\parallel_{p}\$ (11)
-Here, $\mathrm{sg}\left(\cdot\right)$ is the stop-gradient operation preventing the gradient flowing into $G_{e}$ or $G_{d}$, which decouples the feature-pixel domain bidirectional mapping and feature rescaling operations.
-Finally, the loss in the first stage is a combination of these two losses:
-$\mathcal{L}_{\mathrm{stage1}}=\lambda_{1}\mathcal{L}_{\mathrm{rec.}}+\lambda_{2}\mathcal{L}_{\mathrm{gui.}}\$ (12)
+After rescaling, the low-frequency regions of an image typically experience less degradation, while the degradation in high-frequency areas is more conspicuous, necessitating a higher generative capability for restoration.
+Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features.
+We propose to use a simple yet effective network named time-step prediction module (TPM) to regress $t$ based on the rescaled latent features:
+$t=\mathrm{TPM}\left(\hat{z}\right)$ (11)
+Subsequently, the predicted time-step $t$ is used to perform the single denoising step.
+During the training phase, we compute a single time-step for each input image patch.
+In the inference phase, we adopt a tiled inference approach, allowing for predicting different time-steps for various regions of a single image.
+However, the fixed scheduler used in equation (9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM.
+Considering that given $\hat{z}$, $\epsilon$ and $t$, equation (9) essentially represents a linear transformation, which can be learned by a neural network.
+However, if we simply use a randomly initialized neural network to simulate this process, it would cause the output to deviate significantly from the latent space of the pre-trained VAE, leading to training instability.
+Therefore, we propose a hybrid refiner module consisting of both a fixed scheduler and a learnable scheduler.
+The former uses equation (9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler.
+To ensure training stability, the final layer of the learnable scheduler is a zero-initialized convolution, allowing it to gradually correct the results of the fixed scheduler as training progresses.
+This process can be described as follows:
+$\begin{split}\hat{z}_{0}=\mathcal{S}_{\mathrm{fixed}}\left(\hat{z},\epsilon,t_{0}\right)+\mathcal{S}_{\mathrm{learned}}\left(\hat{z},\epsilon,t\right)\end{split}$ (12)
+Here $t_{0}$ represents the fixed pre-set time-step.
+Figure 5: The structure of the refiner module.
+It combines both the fixed scheduler and learnable scheduler to perform the denoising process.
+The two schedulers are connected by a zero-convolution layer to stabilize the training.
 Table 1: Quantitative comparisons with different methods at 16$\times$ and 32$\times$.
@@
 VQIR 23.91 22.02 0.6498 0.5823 0.3174 0.4568 0.1024 0.2663 64.04 58.21 0.6350 0.6293
-Ours 22.58 20.41 0.6171 0.5442 0.3559 0.4763 0.1188 0.1964 67.73 65.27 0.6934 0.6528
+Ours 23.77 22.10 0.6675 0.5839 0.3026 0.4281 0.0824 0.1597 66.31 65.51 0.6443 0.6521
 Urban100 EDSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
@@
 VQIR 20.27 18.44 0.5782 0.4427 0.3038 0.5045 0.1418 0.3290 71.56 61.64 0.6619 0.6568
-Ours 22.16 17.76 0.6809 0.4342 0.2647 0.5164 0.1253 0.2904 72.71 69.72 0.6832 0.6602
+Ours 21.34 18.77 0.6464 0.4655 0.2601 0.4490 0.1107 0.2265 72.21 71.03 0.6773 0.6643
 CLIC2020 EDSR 26.33 23.92 0.7408 0.7062 0.5109 0.5742 0.3057 0.4024 33.17 21.77 0.3392 0.4307
@@
 VQIR 25.69 24.12 0.7363 0.6942 0.2854 0.4191 0.0901 0.2445 64.36 54.83 0.6105 0.6068
-Ours 24.62 22.59 0.7063 0.6551 0.3423 0.4618 0.1217 0.2036 66.48 61.67 0.6662 0.6229
+Ours 25.81 24.16 0.7497 0.6887 0.2777 0.3968 0.0754 0.1459 65.84 64.38 0.6198 0.6586
 DIV8K EDSR 26.56 24.48 0.6776 0.6472 0.5437 0.6048 0.3358 0.4325 30.83 22.30 0.3701 0.4686
@@
 VQIR 25.79 24.26 0.6889 0.6350 0.3221 0.4457 0.1066 0.2628 56.76 51.44 0.5769 0.6203
-Ours 24.39 22.94 0.6515 0.5943 0.3846 0.5058 0.1524 0.2451 58.28 55.68 0.6663 0.6231
-3.3 Perception-oriented Optimization
-The $\hat{z}$ in the first stage can be passed to the decoder $\mathcal{D}$ to obtain the rescaled HR image.
-Although this result has high fidelity, it lacks rich texture.
-Therefore, training in the second stage is designed to enhance perceptual quality.
-Inspired by relevant content generation [32, 19], we adopt the adapter architecture to guide the generation of the SD model.
-The adapter module takes the rescaled features $\hat{z}$ as input and extracts multi-scale features in a coarse-to-fine manner, allowing it to output multi-scale features of different sizes:
-$\{F_{c}^{1},F_{c}^{2},...,F_{c}^{K}\}=\mathrm{Ada}(\hat{z})$ (13)
-Subsequently, these features are element-wisely added to different layers of the denoising U-Net.
-Considering that image rescaling task places a higher emphasis on fidelity, we employ conditional control over both the encoding and decoding layers of the denoising U-Net:
-$F_{\mathrm{enc}}^{i}=F_{\mathrm{enc}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (14)
-$F_{\mathrm{dec}}^{i}=F_{\mathrm{dec}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (15)
-Here, $F_{\mathrm{enc}}^{i}$ and $F_{\mathrm{dec}}^{i}$ represent the output features of the encoder or decoder of the denoising U-Net at the $i$-th layer.
-Finally, the SD model and adapter are optimized using the following loss function:
-$\mathcal{L_{\mathrm{stage2}}}=\mathrm{\mathbb{E}}_{\left(z,\epsilon,\hat{z},t\right)}\parallel\epsilon-\epsilon_{\theta}\left(z_{t},t,\hat{z}\right)\parallel^{2}$ (16)
+Ours 25.93 24.15 0.7107 0.6302 0.3198 0.4376 0.0927 0.1798 57.75 58.89 0.6598 0.6461
 Figure 6: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results.
+The first row represents the original high-resolution images.
 Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
-Figure 6: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both HR images and corresponding edge maps.
+Figure 7: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both the rescaled images and corresponding edge maps.
 Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
+Figure 8: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset
+Figure 9: Visual comparisons of downscaled LR images with sota methods.
+In this part, we fine-tune the denoising U-Net network using LoRA, while also jointly training the TPM and refiner module.
+To achieve a balance between fidelity and realism, we supervise the reconstruction results $\hat{x}$ using pixel loss and perceptual loss as follows:
+$\begin{split}\mathcal{L}_{\mathrm{enh.}}=&\underbrace{\left\|\hat{x}-x\right\|_{1}}_{\text{pixel loss}}+\\ &\underbrace{\lambda_{\mathrm{lpips}}\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\lambda_{\mathrm{dists}}\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)}_{\text{perceptual loss}}\end{split}$ (13)
 4 Experiments
 4.1 Experimental Setup
-4.1.1 Datasets and Evaluation Metrics
-We train the proposed model using the DIV2K dataset [1] and the Flickr2K dataset [17].
-Due to the low resolution of conventional benchmark datasets such as Set5, Set14, and BSD100, they are not suitable for extreme rescaling tasks [5].
-Therefore, we utilized four high-resolution datasets: DIV2K, Urban100 [6], CLIC2020 [23], and DIV8K [5] to evaluate our model.
+4.1.1 Datasets and Evaluation Metrics.
+We train the proposed OSIRDiff using the DF2K dataset [1, 19].
+Considering that traditional benchmarks such as Set5, Set14, and BSD100 often have lower quality and resolution [5], they are not suitable for evaluating extreme rescaling methods.
+Therefore, we utilize four high-resolution datasets: DIV2K, Urban100 [7], CLIC2020 [32], and DIV8K [5] to evaluate models been compared.
 We adopt PSNR and SSIM [37] for quantitative evaluation.
@@
 4.1.2 Training Details
-Our LSDIR is built based on the SD 2.1-base model and the training is divided into two stages.
-First, we train the DFRM together with the pre-trained VAE model for 200K iterations using Eq.
-12 The learning rate is initialized as $2\times 10^{-4}$ and reduced by half every 50k iterations.
-After the training of DFRM, we adopt Eq.
-(16) to train the Adapter and finetune the diffusion U-Net.
-We only finetune the linear layers of the U-Net for its better performance.
-During both stages, we adopt Adam optimizer [12] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and set the mini-batch size to 32.
+OSIRDiff is built based on the SD 2.1-base model, and the training is divided into three stages.
+First, we train the DFRM using equation (8).
+Then we adopt equation (13) to jointly train the LoRA layers, the TPM and the refiner module.
+Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and refiner module using a combination of equations (8) and (13).
+More training details are provided in the supplementary material.
 4.2 Results and Comparison with SOTA
+We compare the proposed OSIRDiff with three kinds of methods: (1) bicubic downscaling and SR, i.e., EDSR [19], RCAN [50] and ESRGAN [36]; (2) flow-based rescaling models, i.e., IRN [41] and HCFlow [18]; (3) prior-based extreme rescaling method VQIR [39].
+Some of them are proposed for 4$\times$ rescaling factors and we extended them to 16$\times$ or 32$\times$.
+To ensure a fair comparison, all methods are trained until convergence on the DF2K dataset following their original configurations.
+4.2.1 Quantitative Results
 […]
-4.2.1 Quantitative Results
-Tab.
-1 presents the 16$\times$ and 32$\times$ quantitative image rescaling results on four datasets.
-The performance of SR methods are generally poor.
-While methods based on INNs achieve the highest PSNR and SSIM scores due to consideration of the correlation between downscaling and upscaling.
-VQIR exhibits the best LPIPS by fully utilizing the prior provided by pre-trained VQGAN and being trained based on LPIPS-loss.
-As our proposed method is trained in the latent space, the results exhibit the best semantic integrity, which leads to significant improvements in two non-reference metrics.
-It is important to note that while our method slightly lags behind VQIR in terms of certain full-reference metrics, it excels in reconstructing accurate semantics and regular structures, as will be demonstrated in the following section.
 4.2.2 Qualitative Results
 In Fig.
-5, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors.
+6, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors.
 Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as EDSR, RCAN, IRN, and HCFlow produce overly-smoothed results.
@@
 It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as faces, text and architectural structures.
-The quantitative experiments demonstrate that our model exhibits outstanding performance on the Urban100 dataset, confirming that our method produces more accurate structural features.
-Therefore, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig.
-6.
-Traditional regression models, such as EDSR, RCAN, IRN, and HCFlow, tend to generate overly-smooth results with few sharp edges.
-Although GAN-based methods, including ESRGAN and VQIR, generate more details, their edge information lacks accuracy and regularity.
-In contrast, our approach, by leveraging the prior knowledge encapsulated in the Stable Diffusion model, is able to reconstruct more accurate and rich edge details.
+The quantitative experiments demonstrate that our model exhibits superior performance on the Urban100 [7] dataset, confirming that our method produces more accurate structural features.
+In this section, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig.
+7.
+Traditional regression models, such as EDSR [19], RCAN [50], IRN [41], and HCFlow [18], tend to generate overly-smooth results.
+GAN-based methods, including ESRGAN [36] and VQIR [39], generate more details, but their edge information lacks accuracy and regularity.
+In contrast, our approach, by leveraging the prior knowledge encapsulated in the SD model, is able to reconstruct more accurate and rich edge details.
+In Fig.
+8, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K [5].
+It is evident that our method is capable of recovering images with accurate semantics, even at extreme upscaling factors.
+To validate the effectiveness of our downscaling scheme, we conduct a visual comparison of the downscaled LR images with state-of-the-art (SOTA) methods, as shown in Fig.
+9.
+Overall, the LR images generated by our method achieve comparable or even superior visual quality to SOTA image rescaling methods.
+The LR images produced by VQIR exhibit noticeable color shifts.
+Furthermore, VQIR generated LR images contain significant noise, whereas our method exhibits a certain degree of ringing artifacts.
+In future work, we plan to further optimize the quality of the LR images.
+Figure 10: The validation loss curves when removing a specific component from the DFRM.
 4.3 Ablation Study
@@
 To verify the effectiveness of the feature rescaling operation, we move the rescaling operation from the latent space to the pixel space.
-This means we directly train a downsampling network to downscale the HR $x$ to LR $y$.
-Subsequently, $y$ is used as a condition to guide the generation process of the SD model.
-As shown the 1st row in Tab.
-2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by lower LPIPS and DISTS.
-In Fig.
-7, we plot the LPIPS variation curves on the validation set for the two rescaling schemes.
-The pixel space-based rescaling scheme requires a longer time to converge, whereas the latent space based scheme can stably converge to a better result in a shorter time.
-Additionally, we observe that the pixel space based scheme experiences severe color shifts in the middle of training, as illustrated in Fig.
-8.
-It requires a substantial number of iterations to learn the global color distribution of the reconstructed images.
-This is because using pixels as the conditioning input for Stable Diffusion makes it difficult to align with the intrinsic priors of the pre-trained model, leading to issues such as unstable training and longer training time.
-Figure 7: The validation LPIPS curve for rescaling models based on latent space and pixel space.
-Figure 8: Reconstructed results after 60k training steps. The pixel space based scheme shows severe color shifts.
+Specifically, we use HCFlow [18], the comparative method with the best PSNR performance, to perform the rescaling operation in the pixel domain, followed by perceptual enhancement using the same approach described in Sec.
+3.3.
+As shown in Fig.
+11, the rescaling operator in pixel space takes HR image as input and simultaneously outputs both the LR image and the rescaled image.
+Then, to use SD for perceptual enhancement, we need to map the rescaled image to the latent space using a VAE encoder and perform the denoising process.
+Finally, the enhanced latent features are mapped back to the pixel space using the VAE decoder.
+Due to the nonlinear mapping nature of the VAE encoder, the minimum distance in pixel space does not correspond to the minimum distance in latent space.
+This causes the rescaling operator trained in pixel space to misalign with the prior of the pre-trained SD model, resulting in a degradation of perceptual quality.
+As shown the 1st row in Table 2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by much lower LPIPS and DISTS.
+Figure 11: Image rescaling within pixel space and latent space with Stable Diffusion (SD) prior.
+As shown in Fig.
+12, we conduct visual comparisons between rescaling in latent space and pixel space in two different scenarios.
+Our latent-space rescaling method is able to reconstruct more realistic textures and more accurate structural features.
+This indicates that latent-space rescaling operation can preserve sufficiently fine-grained contextual information about the HR image, such as structure, texture, and semantics.
+Figure 12: Visual comparisons of our method with rescaling in pixel space and without the SD prior.
 Table 2: Ablation study of rescaling space and SD prior.
 Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
-$\times$ $\checkmark$ $\checkmark$ 0.3818 0.1333
+$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
 $\checkmark$ $\times$ $\times$ 0.4675 0.3109
-$\checkmark$ $\times$ $\checkmark$ 0.3559 0.1188
-4.3.2 Effectiveness of SD Prior
-To validate that SD model provides rich priors about real-world HR images, we directly omit the second training stage.
-As shown in Tab.
-2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results.
+$\checkmark$ $\times$ $\checkmark$ 0.3026 0.0824
+4.3.2 Necessity of SD Prior
+To validate that SD model provides rich priors about real-world HR images, we directly omit the perceptual enhancement part.
+As shown in Table 2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results.
+Here, we compare the latent features before and after perceptual enhancement by decoding them to the pixel space using the VAE decoder, as shown in Fig.
+12.
+It can be seen that the latent features before perceptual optimization exhibit noticeable color shifts and blurriness when mapped to the pixel domain.
+In contrast, the perceptually enhanced latent features display richer textures and more accurate color information.
+Therefore, leveraging the rich natural image priors in the pre-trained Stable Diffusion model can significantly enhance the visual quality of rescaled images.
 4.3.3 Effectiveness of DFRM Design
-Our DFRM includes three specific designs: the domain converter module, the pixel guidance module, and the stop-gradient operation in the loss function Equ.
-11.
+Our DFRM includes two specific designs: the domain converter module and the pixel guidance module.
 To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig.
 10.
-It can be observed that the introduction of the INN and the stop-gradient operation can both reduce reconstruction error.
-Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as shown in Tab.
-3.
+It can be observed that the introduction of the INN can reduce reconstruction error.
+Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as demonstrated by the lower guidance loss in Fig.
+10.
+As shown in Fig.
+13, the LR images exhibit significant noise and color distortions in the absence of pixel guidance.
+However, the introduction of pixel guidance significantly alleviates these issues.
 Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
-Ideally, the domain converter composed of a set of invertible neural networks should be able to accurately perform bidirectional mapping between the feature domain and the pixel domain, enabling the LR to share the same content with HR images.
-However, due to the limited representation power of INN, the intermediate features $w$ obtained by simply downscaling the latent code $z$ are not well-suited for conversion to the pixel domain.
-The introduction of pixel guidance allows for embedding some pixel-domain information into the intermediate features $w$, thereby achieving a balance between reconstruction performance and LR image quality.
+Figure 13: Visual comparisons of LR images with and without pixel guidance.
+Figure 14: Ablation study about time alignment.
+4.3.4 Effectiveness of Time Alignment
+To validate the effectiveness of our proposed time-step alignment strategy, we retrain our model using fixed time-steps.
+This involves removing the TPM and refiner modules and employing a manually selected time-step to perform a single denoising step based on equation (9).
+Following the setting of previous work [40], we select fixed time-steps 1 and 999 for comparison.
 As shown in Fig.
-9, the LR images exhibit significant noise and color distortions in the absence of pixel guidance.
-However, the introduction of pixel guidance significantly alleviates these issues.
-Figure 9: Comparisons of LR images visual quality with and without pixel guidance.
-Table 3: Ablation study of pixel guidance module. The introduction of pixel guidance can significantly enhance the quality of LR images.
-Method PSNR\ $\uparrow$ SSIM\ $\uparrow$ NIQE\ $\downarrow$
-w/o pixel guidance 19.88 0.5898 22.33
-with pixel guidance 21.99 0.7237 13.27
-Figure 10: The validation loss curves when removing a specific component from the DFRM.
-4.4 Discussions
-4.4.1 Comparison with JPEG Compression
-Our model can be combined with existing image compression methods to achieve more extreme compression ratios.
-To validate this, we use the lossless compression method PNG to compress LR images and compare its performance with JPEG compression at various LR sizes.
-Specifically, we train rescaling models at three different scales (8$\times$, 16$\times$, and 32$\times$) and record the bpp by the LR images.
-We also adjust the quality parameters of JPEG compression accordingly for comparison.
+14, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen.
+In contrast, our time-step alignment mechanism can dynamically determine the level of generative capability based on the image content, thereby achieving optimal performance in both fidelity and perceptual quality.
+5 Discussions
+5.1 Comparison with JPEG Compression
+Our model can be combined with existing lossless image compression methods to achieve extreme compression ratios.
+To validate this, we use the lossless compression image format PNG to compress LR images and compare its performance with JPEG compression at various bit-rates.
 As shown in Fig.
-11, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of JPEG compression.
+15, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression.
+These comparisons demonstrate the potential of our model for image compression.
+Figure 15: R-D curves of JPEG compression and our proposed model on the DIV2K validation set.
 In Fig.
-12, we present the visual quality, the bpp and the LPIPS of the reconstructed images.
+16, we present the visual quality, the bpp and the LPIPS of the reconstructed images.
 Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics.
 Additionally, JPEG compression tends to produce noticeable block artifacts and color distortions in the background regions of the images, whereas our model maintains superior reconstruction quality across all three scales.
-These comparisons demonstrate the potential of our model for image compression.
-Figure 11: R-D curves of JPEG compression and our proposed model on the DIV2K validation set.
-Figure 12: Visual comparisons between our model and JPEG compression. Below the images, the bpp and LPIPS of the reconstructed images are provided.
-4.4.2 Fidelity-realism Trade-off
-Our two-stage model outputs two features in the latent space ($\hat{z}$ and $\hat{z}_{0}$), which represent reconstruction results with high fidelity and high realism respectively.
-It is important to note that the high fidelity mentioned here does not refer to higher pixel fidelity but to semantic fidelity in the latent space.
-By fusing the results of the two stages, we can achieve a trade-off between fidelity and realism:
-$z_{\mathrm{fus.}}=\alpha\hat{z}+\left(1-\alpha\right)\hat{z}_{0}$ (17)
-$x_{\mathrm{fus.}}=\mathcal{D}\left(z_{\mathrm{fus.}}\right)$ (18)
-As shown in Fig.
-13, by adjusting the value of $\alpha$, we can achieve a smooth transition between fidelity and realism.
-More results are provided in the supplementary materials.
-As shown in Fig.
-14, It can be observed that by adjusting the fusion weights, we can achieve a smooth transition between fidelity and realism.
-Figure 13: By adjusting the fusion ratio of two latent codes, we can achieve a trade-off between fidelity and realism.
-Figure 14: By adjusting the fusion weights, our model can achieve a trade-off between fidelity and realism.
-The perceptually optimized results display richer details, but they also exhibit some artifacts.
-Figure 15: Visual comparisons of LR images.
+Figure 16: Visual comparisons between our model and JPEG compression. bpp/LPIPS of the reconstructed images are shown below the images.
+5.2 Time-step Distribution
+To overcome the non-uniform degradation introduced by the rescaling operation, we design a time-step alignment strategy to predict an optimal time-step $t$ based on the degraded latent variable $\hat{z}$, enabling dynamic allocation of generative capability.
+During the inference phase, our model can perform tiled inference, allowing it to predict different time-steps for various regions of the same image.
+We discuss the distribution of time-steps at both the single image and dataset scales.
+Figure 17: The time-steps $t$ predicted by the TPM on a single image, along with the corresponding latent variable rescaling error map $|z-\hat{z}|$.
+First, we visualize the time-steps corresponding to different regions of the image shown in Fig.
+17.
+We also present the latent variable rescaling error map $|z-\hat{z}|$ for comparison.
+It can be observed that in more complex regions, the rescaling operation introduces larger errors.
+Correspondingly, these areas also require larger time-steps.
+This is because complex regions necessitate greater generative capability for restoration.
+Figure 18: The distribution of time-steps predicted by the TPM on the DIV2K validation set at different rescaling scale factors. Smaller scale factor results in smaller time-steps.
+Next, we present the distribution of predicted time-steps at different rescaling scales on the DIV2K validation set shown in Fig.
+18.
+The vertical axis and the size of the bubbles represent the mean and variance of the time-steps, respectively.
+It can be observed that the mean of the predicted time-steps increases as the rescaling scale becomes larger.
+This is because larger rescaling scales result in greater image degradation, necessitating stronger generative capability for restoration.
+Additionally, we find that the predicted time-steps for the 16$\times$ rescaling exhibit the largest variance.
+This indicates that 16$\times$ rescaling is a task that balances both fidelity and generative capacity.
+In contrast, 8$\times$ rescaling tends to prioritize fidelity, while 32$\times$ rescaling leans more towards generative capabilities.
 6 Limitations
-Although our model demonstrates superior performance in extreme image rescaling tasks, it still has some limitations.
-One of them is the quality of the LR images.
-As shown in Fig.
-15, the LR images generated by IRN closely resemble the bicubic downsampled image.
-In contrast, the LR images produced by HCFlow and VQIR exhibit some color shifts.
-LR images generated by our model are slightly noisy.
-In the future, more advanced invertible network architectures and more effective loss functions could be explored to further improve the quality of LR images.
+Table 3: The performance comparison between our 4×4$\times$4 × rescaling model and VAE.
+Method PSNR $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
+VAE 29.00 0.1570 65.54 0.6234
+Ours-4×4$\times$4 × 29.32 0.1456 65.47 0.6313
+6.1 Upper Bound of OSIRDiff
+Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance ceiling is inherently constrained by the VAE itself.
+The VAE encoding is a lossy process and can be described as follows:
+$z=\mathcal{E}(x)$ (14)
+Here, $x\in R^{H\times W\times 3}$ represents original HR image and $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ represents corresponding latent features.
+We can measure the compression rate from the perspective of the number of bits:
+[…]
+$=\left(H\times W\times 3\right)\times 8$ (15)
+[…]
+$=\left(\frac{H}{8}\times\frac{W}{8}\times 4\right)\times 16$ (16)
+$\mathrm{Rate}$ $=\frac{\mathrm{Bits(z)}}{\mathrm{Bits(x)}}=\frac{1}{24}$ (17)
+It can be observed that VAE encoding can be regarded as a compression process with a compression ratio of 24.
+From the perspective of rescaling, this is equivalent to a rescaling operation with a scale factor of $S=\sqrt{24}$.
+Any rescaling scale smaller than $S$ will inevitably be limited by the reconstruction performance of VAE, as the VAE encoding process has already discarded significant amounts of information.
+Furthermore, we directly encode and decode input HR images using VAE without performing any rescaling operations, representing the upper performance bound of our model.
+Additionally, we train an OSIRDiff model for 4×4$\times$4 × rescaling and compare its performance with the reconstruction capability of the VAE, as illustrated in Table 3.
+It can be observed that the performance of our 4×4$\times$4 × rescaling model is very close to the reconstruction performance of the VAE, indicating that the rescaling performance at this scale is limited by the VAE reconstruction capacity.
+This also highlights a potential limitation of SD prior itself that it only demonstrates significant advantages when handling highly ill-posed tasks [8, 2, 46].
+Consequently, our proposed OSIRDiff is similarly affected by this limitation and performs well primarily at larger rescaling scales (>$S$).
+In the future, we could explore the incorporation of shallow features into the VAE decoding process to enhance its fidelity, thereby enabling the OSIRDiff framework to be more applicable to small-scale rescaling tasks.
+Table 4: Comparison of inference time and parameter sizes across methods.
+Time (s) Total Parameters (M) Trainable Parameters (M)
+IRN 0.359 34.2 34.2
+HCFlow 0.363 7.24 7.24
+VQIR 0.897 92.9 20.7
+Ours 0.711 1352.5 63.2
+6.2 Inference Time and Model Size
+In Table 4, we compare the inference speed and parameter sizes of different methods.
+All tests are performed on an RTX 3090 using images with a resolution of $1024\times 1024$.
+Due to leveraging the SD architecture, our method has the largest total number of parameters and the highest number of trainable parameters.
+However, since our approach only requires a single denoising step, its inference time does not increase significantly compared to other methods.
+The inference speed of our method even surpasses that of the VQGAN-based [4] method, i.e.
+the VQIR.
+This is because, compared to VAE, the latent space of VQGAN has a higher number of channels.
+Additionally, the nearest-neighbor search process in VQGAN incurs significant computational overhead, adversely affecting its inference speed.
+In the future, we plan to explore architecture distillation techniques to further reduce the size of our image rescaling model and improve its inference speed.
 7 Conclusion
-Extreme image rescaling is a highly challenging yet significant task.
-In this paper, we propose a novel framework called LSBIR to address issues such as insufficient texture and semantic inaccuracies.
-Specifically, we leverage the rich priors encapsulated in the pre-trained SD model to mitigate the ill-posed nature of rescaling.
-In the first stage, we design a DFRM to achieve bidirectional mapping between HR features and LR images.
-In the second stage, the rescaled features from the first stage serve as conditions to guide the generation process of the SD model through an effective adapter.
+We propose a novel framework called OSIRDiff to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling.
+Specifically, we design a DFRM to achieve bidirectional mapping between HR features and LR images.
+Then, we utilize a pre-trained SD model to perform a single denoising step on the rescaled features, enhancing its perceptual quality.
+Considering the non-uniform quality of the rescaled latent features, we propose a novel time-step alignment strategy to achieve a balance between fidelity and perceptual quality.
 Both qualitative and quantitative experiments demonstrate the superiority of our approach.

### v3

@@
-One-step Generative Diffusion for Realistic Extreme Image Rescaling
+Timestep-Aware Diffusion Model for Extreme Image Rescaling
 Abstract
@@
 However, extreme downscaling factors pose significant challenges to the upscaling process due to its highly ill-posed nature, causing existing image rescaling methods to struggle in generating semantically correct structures and perceptual friendly textures.
-In this work, we propose a novel framework called One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model.
-Specifically, OSIRDiff adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image.
-Then, the rescaled features are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details.
-The entire model is end-to-end trained to enable the diffusion priors to guide the rescaling process.
-Considering the spatially non-uniform reconstruction quality of the rescaled latent features, we propose a novel time-step alignment strategy, which can adaptively determine the generative strength of the diffusion model based on the degree of latent reconstruction errors.
-Extensive experiments demonstrate the superiority of OSIRDiff over previous methods in both quantitative and qualitative evaluations.
+In this work, we propose a novel framework called Timestep-Aware Diffusion Model (TADM) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model.
+Specifically, TADM adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image.
+Then, the rescaled latent features are enhanced by a pre-trained diffusion model to generate more faithful details.
+Considering the spatially non-uniform degradation caused by the rescaling operation, we propose a novel time-step alignment strategy, which can adaptively allocate the generative capacity of the diffusion model based on the quality of the reconstructed latent features.
+Extensive experiments demonstrate the superiority of TADM over previous methods in both quantitative and qualitative evaluations.
 The code will be available at: https://github.com/xxx/xxx.
@@
 When HR images are needed, an upscaling process is employed to restore the downscaled images to their original resolution.
-Unlike image super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
-Figure 1: Visual comparison of 16$\times$ image rescaling results.
-Recently published image rescaling methods can be classified into two main categories: encoder-decoder architecture [14, 31] and invertible neural networks [41, 18].
+Unlike image super-resolution (SR), image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
+Recently proposed image rescaling methods can be classified into two main categories: encoder-decoder architecture [12, 24] and invertible neural networks [36, 17].
 The encoder-decoder architecture employs two neural networks to perform downscaling and upscaling.
@@
 Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results.
-GRAIN [53] and BDFlow [15] use pre-trained StyleGAN [10, 11] to generate high-quality upscaling face images.
+GRAIN [46] and BDFlow [14] use pre-trained StyleGAN [9, 10] to generate high-quality upscaled face images.
 VQIR [33] leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [3] to achieve extreme rescaling on natural images.
-Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from incorrect semantics and structure.
-As shown in Fig.
-1, although VQIR generates richer details compared to HCFlow, it still fails to reconstruct accurate text and image patterns.
-To address the aforementioned issues, we propose a novel framework named One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images.
-Specifically, first, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image.
-Then, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality.
-Finally, the entire model is trained end-to-end to allow the diffusion priors to guide the rescaling process.
-Due to the spatially non-uniform reconstruction quality of the rescaled latent features, different regions of the image require varying levels of generative capability for restoration.
+Although these methods have made significant progress in perceptual metrics, their reconstruction results still suffer from incorrect semantics and structure.
+To address the aforementioned issues, we propose a novel framework named Timestep-Aware Diffusion Model (TADM) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images.
+Specifically, first, the HR images are encoded by the via the VAE encoder to obtain the latent features.
+Then, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image.
+Next, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality.
+Finally, the enhanced features are decoded via the VAE decoder to obtain the rescaled image.
+Additionally, we find that the rescaled latent features exhibit non-uniform degradation, which can be interpreted as the addition of noise at varying levels.
+As a result, different images require different time-steps to perform the denoising process.
 Previous studies have shown that the time-step in diffusion models directly controls the generative capacity of the model [28].
-Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby achieving a balance between fidelity and realism.
-As shown in Fig.
-1, our model significantly enhances the overall quality of the rescaled images compared to other methods.
+Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby adaptively tuning the generative capacity of the diffusion model.
 Our contributions can be summarized as follows:
-  • We propose a One-Step Image Rescaling Diffusion framework that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
+  • We propose a Timestep-Aware Diffusion Model that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
   • We propose a novel decoupled strategy for feature rescaling.
@@
   • We design a time-step alignment strategy to address the spatially non-uniform reconstruction quality of the rescaled latent features, allowing for more effective utilization of the prior information from the text-to-image diffusion model.
-Figure 2: Overview of the proposed One-Step Image Rescaling Diffusion (OSIRDiff).
+Figure 1: Overview of the proposed Timestep-Aware Diffusion Model (TADM).
 First, the HR images $x$ are encoded to obtain the latent features $z$.
-These latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$.
+The latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$.
 Next, the rescaled latent features $\hat{z}$ are passed into a pre-trained diffusion model to perform a single denoising step for perceptual enhancement to obtain $\hat{z}_{0}$.
+In this process, a time prediction module (TPM) is used to estimate the time step $t$ based on $\hat{z}$ and the predicted $t$ is then fed into both the U-Net and the time scheduler.
 Finally, the perceptually enhanced latent features $\hat{z}_{0}$ are decoded to obtain the rescaled image $\hat{x}$.
@@
 Nonetheless, these approaches are often limited in image contents, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
-Recently, the emergence of the pre-trained Stable Diffusion (SD) model [27] has showcased its efficacy in both content generation [47] and image restoration [34].
-Despite this progress, to date, no effort has been made to investigate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
 2.2 Diffusion Model for Image SR
 Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR.
-Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model.
+Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch [13, 4, 40] and leveraging pre-trained SD model [19, 39, 35].
 The former is represented by SR3 [23], which achieves perceptual quality comparable to GANs.
 The latter is represented by StableSR [27], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios.
-However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling.
-In this work, we decompose extreme image rescaling into two sub-tasks: feature rescaling in the latent space and one-step perception enhancement with the pre-trained SD U-Net.
-2.3 Diffusion Acceleration
-To accelerate the inference process of diffusion models, researchers have proposed various optimization strategies.
-Fast diffusion samplers [30, 12, 21, 51] effectively reduce the sampling steps required by pre-trained diffusion models, lowering them from the initial thousand steps to as few as 20–50 steps.
-However, further reductions often lead to substantial performance degradation.
-Distillation techniques [23, 29, 42, 52, 45, 24] have shown significant promise in accelerating inference.
-For example, Progressive Distillation [29] minimizes sampling steps to under 10 by training a sequence of student models with progressively increasing stride.
-Instaflow [20] advances this further, achieving one-step generation using reflow and distillation methods.
-Recently, score-based distillation approaches, including Distribution Matching Distillation [45] and Variational Score Distillation [24], have utilized pre-trained text-to-image diffusion models as distribution matching losses to enable one-step text-to-image generation.
-Although these methods have made significant strides, high training costs remain a challenge.
-Inspired by Diffusion2GAN [9], we propose a low-overhead training approach based on Low-Rank Adaptation (LoRA) [6], achieving one-step acceleration for diffusion models and adaptability for image rescaling tasks.
+However, the above methods typically require dozens or even hundreds of iterations during inference, which significantly increases computational costs.
+Researchers have begun exploring how to achieve image SR using one-step diffusion models.
+SinSR [30], based on ResShift [40], derives a deterministic one-step sampling process and employs a consistency preserving loss to distill the prior knowledge from a multi-step teacher network into a one-step student network.
+A representative work, OSEDiff [34], directly feeds the LQ image into the SD model and employs variational score distillation for regularization.
+S3Diff [42] proposes a degradation-aware LoRA and utilizes a pre-trained DINO [43] model as a discriminator for adversarial training.
+InvSR [41] designs a partial noise prediction strategy to provide a dynamic starting sampling point and supports inference with an arbitrary number of sampling steps ranging from one to five.
 3 Methodology
-3.1 Overview of OSIRDiff
+3.1 Overview of TADM
 As illustrated in Fig.
-2, OSIRDiff consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, One-Step Perception Enhancement and Latent Decoding.
+1, TADM consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, Denoising Guided Perception Enhancement and Latent Decoding.
 First, the input HR images $x$ are mapped to the latent features $z$ via the pre-trained VAE encoder $\mathcal{E}$.
 Subsequently, the proposed Decoupled Feature Rescaling Module (DFRM) rescales $z$ to the target size, outputting LR images $y$ and the rescaled latent features $\hat{z}$.
-Our aim is to make $\hat{z}$ as close as possible to $z$ so that the rescaled results align with the latent space of the pre-trained VAE, while ensuring that $y$ and the original HR images $x$ share the same content.
 Then, the rescaled features $\hat{z}$ are fed into a denoising U-Net network to perform a single denoising step to get the perceptually enhanced latent features $\hat{z}_{0}$.
-Finally, perceptually enhanced latent features $\hat{z}_{0}$ are decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$.
-To dynamically adjust the generative capability based on the reconstruction quality of the rescaled latent features, we design a time-step alignment strategy, which consists of a time-step prediction module (TPM) and a refiner module.
+Finally, $\hat{z}_{0}$ is decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$.
+Overall, the latent rescaling process can be regarded as a noise-adding process, while the latent enhancement process can be viewed as a denoising process.
+To dynamically estimate the ‘noise’ density introduced by latent rescaling, we design a time alignment strategy, which consists of a time-step prediction module (TPM) and a time scheduler module.
+It can better address the non-uniform degradation caused by rescaling, thereby enabling the dynamic allocation of generative capabilities of pre-trained SD.
 3.2 Feature Rescaling in the Latent Space
@@
 In this part, we perform rescaling of the HR latent features $z$, obtaining LR images $y$ and rescaled features $\hat{z}$.
-A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$.
-Here, $G_{e}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{d}$ represents the decoder used to reconstruct the original HR features.
-Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
+A straightforward idea is to construct a transformation chain as $y=G_{\mathrm{e}}\left(z\right)$, $\hat{z}=G_{\mathrm{d}}\left(y\right)$.
+Here, $G_{\mathrm{e}}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{\mathrm{d}}$ represents the decoder used to reconstruct the original HR features.
+Guidance loss $\mathcal{L}_{\mathrm{gui.}}$ and reconstruction loss $\mathcal{L}_{\mathrm{rec.}}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
 $\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{1}$ (1)
@@
 However, these two losses are contradictory to each other.
-When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [38], and little information regarding the downscaling process can be embedded into the LR image.
-In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [22], which results in the LR image lacking any visual content.
-Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ could adversely impact subsequent upscaling.
+When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [32], embedding little information about the downscaling process into the LR image.
+In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [20], causing the LR image to lack meaningful visual content.
+Therefore, directly applying the guidance loss $\mathcal{L}_{\mathrm{gui.}}$ on $y$ could adversely impact subsequent upscaling.
 Figure 2: Detailed architecture of Decoupled Feature Rescaling Module (DFRM).
 Considering the above issues, we devise the DFRM as shown in Fig.
-3, which divides the rescaling operation into two sub-tasks: feature rescaling and LR image generation.
+2, which divides the rescaling operation into two sub-tasks: feature rescaling and LR image generation.
 Inspired by FGRN [15], we use a set of CNNs to perform downscaling and upscaling, and a separate INN to perform the bidirectional mapping between the feature domain and the pixel domain.
@@
 Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}\right)$; $\left(z_{\text{lr}}\leftrightarrow y\right)$ as follows:
-$z_{\text{lr}}=G_{e}\left(x,z\right)$ (3)
+$z_{\text{lr}}=G_{\mathrm{e}}\left(x,z\right)$ (3)
 $z_{\text{lr}}\approx F^{-1}\left(F\left(z_{\text{lr}}\right)\right)=F^{-1}\left(y\right)$ (4)
-$\hat{z}=G_{d}\left(z_{\text{lr}}\right)$ (5)
+$\hat{z}=G_{\mathrm{d}}\left(z_{\text{lr}}\right)$ (5)
 Here, $F$ and $F^{-1}$ represent the forward and backward processes of the INN, respectively.
-Given the latent variable $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $z_{\text{lr}}\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor.
-Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization.
-Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$.
-Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
+Given the latent variable $z\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{\mathrm{e}}$ first transforms it into a compact representation $z_{\text{lr}}\in\mathbb{R}^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor.
+Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in\mathbb{R}^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization.
+In the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$.
+Finally, the decoder $G_{\mathrm{d}}$ is utilized to obtain the rescaled feature $\hat{z}$.
 3.2.2 Loss Function.
@@
 Considering the information lost in quantization operations, the bidirectional mapping between $z_{\text{lr}}$ and $y$ is not exactly reversible.
-To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{y}_{lr}\rightarrow\hat{z}$, which can be expressed as the following formula:
-$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
+To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{z}_{\text{lr}}\rightarrow\hat{z}$, which can be expressed as the following formula:
+$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{\mathrm{d}}\left(G_{\mathrm{e}}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{\mathrm{d}}\left(F^{-1}\left(F\left(G_{\mathrm{e}}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
 This combination can facilitate the network to reconstruct $\hat{z}$ with high fidelity in both the existence and absence of quantization operations.
 Additionally, the guidance loss is designed as follows:
-$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{e}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
+$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{\mathrm{e}}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
 Finally, the loss function of this part is a combination of these two losses:
 $\mathcal{L}_{\mathrm{res.}}=\lambda_{\mathrm{rec.}}\mathcal{L}_{\mathrm{rec.}}+\lambda_{\mathrm{gui.}}\mathcal{L}_{\mathrm{gui.}}\$ (8)
-Figure 4: The rescaling operation leads to spatially non-uniform reconstruction quality of the rescaled latent features, which depends on the content of the image and the rescaling factor.
-3.3 One-Step Perception Enhancement
+3.3 Denoising Guided Perception Enhancement
 The latent variable $\hat{z}$ obtained in Sec.
 3.2 retains sufficient semantic information about the HR image, but lacks adequate texture.
-Therefore, we are proposed to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part.
-Considering the original diffusion model requires a multi-step iterative sampling algorithm, this significantly increases inference time and makes it difficult to scale to practical applications.
+Therefore, we propose to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part.
+Original diffusion model requires a multi-step sampling algorithm, significantly increases inference time and makes it difficult to scale to practical applications.
 Inspired by recent work on accelerating diffusion models [8], we adopt an one-step inference strategy.
-Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ to perform a single denoising step as follows:
+Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ and perform a single denoising step as follows:
 $\hat{z}_{0}=\frac{\hat{z}-\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}$ (9)
 where $\hat{z}_{0}$ represents the perceptually enhanced latent features, $\bar{\alpha}_{t}$ is the noise scheduler of the forward diffusion process of the SD model, and $t$ represents the time-step associated with the noise intensity.
-Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}$:
-$\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$ (10)
+Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$.
 Then, we can adopt the loss function between $\hat{x}$ and $x$ to optimize the U-Net using LoRA layers.
-It should be noted that one of the key designs is the selection of the time-step $t$.
-Previous studies [35] have shown that the time-step in diffusion models could determine the trade-off between realism and fidelity.
-A larger $t$ results in stronger diversity and greater realism, but at the cost of fidelity.
-However, we have observed that the quality of the rescaled latent features exhibit non-uniform characteristics, as shown in Fig.
+However, the key problem of the aforementioned approaches lies in the selection of time-step $t$.
+We can regard the rescaling process as the noise-adding process of diffusion model, while Eq.
+(9) represents the denoising process.
+For different rescaling factors and different images, the ‘noise’ introduced by the rescaling operation varies in intensity.
+In Fig.
+3, we present the global rescaling error of an image under different rescaling factors, as well as the local rescaling error of different patches within the image.
+Additionally, we plot the error curve of the forward noise-adding process of the SD model.
+It can be observed that as the rescaling factor and image content change, rescaling MSE also varies and corresponds to the diffusion MSE at different time-steps.
+Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features $\hat{z}$.
+A simple yet effective network named time-step prediction module (TPM), whose structure is presented in Fig.
+1, is used to predict $t$ based on the rescaled latent features: $t=\mathrm{TPM}\left(\hat{z}\right)$.
+During the training phase, we compute a single time-step for each input image patch.
+In the inference phase, we adopt a tiled inference approach (which will be elaborated in detail in the supplementary materials), allowing for predicting different time-steps for various regions of a high-resolution input image.
+Figure 3: We compare the MSE introduced by the rescaling operation with that caused by the forward diffusion process.
+The rescaling degradation severity is correlated with both the image content and the rescaling scale.
+Therefore, the model requires adaptive time-step to align the rescaling MSE with the diffusion MSE.
+Subsequently, the predicted time-step $t$ is used to perform a single denoising step.
+However, the vanilla SD time scheduler as shown in Eq.
+(9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM.
+Considering that given $\hat{z}$, $\epsilon$ and $t$, Eq.
+(9) essentially represents a linear transformation, which can be learned by a neural network.
+Therefore, we propose a hybrid time scheduler module consisting of a fixed scheduler and a learnable scheduler, whose structure is presented in Fig.
 4.
-After rescaling, the low-frequency regions of an image typically experience less degradation, while the degradation in high-frequency areas is more conspicuous, necessitating a higher generative capability for restoration.
-Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features.
-We propose to use a simple yet effective network named time-step prediction module (TPM) to regress $t$ based on the rescaled latent features:
-$t=\mathrm{TPM}\left(\hat{z}\right)$ (11)
-Subsequently, the predicted time-step $t$ is used to perform the single denoising step.
-During the training phase, we compute a single time-step for each input image patch.
-In the inference phase, we adopt a tiled inference approach, allowing for predicting different time-steps for various regions of a single image.
-However, the fixed scheduler used in equation (9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM.
-Considering that given $\hat{z}$, $\epsilon$ and $t$, equation (9) essentially represents a linear transformation, which can be learned by a neural network.
-However, if we simply use a randomly initialized neural network to simulate this process, it would cause the output to deviate significantly from the latent space of the pre-trained VAE, leading to training instability.
-Therefore, we propose a hybrid refiner module consisting of both a fixed scheduler and a learnable scheduler.
-The former uses equation (9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler.
+The former uses Eq.
+(9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler.
 To ensure training stability, the final layer of the learnable scheduler is a zero-initialized convolution, allowing it to gradually correct the results of the fixed scheduler as training progresses.
@@
 $\begin{split}\hat{z}_{0}=\mathcal{S}_{\mathrm{fixed}}\left(\hat{z},\epsilon,t_{0}\right)+\mathcal{S}_{\mathrm{learned}}\left(\hat{z},\epsilon,t\right)\end{split}$ (10)
-Here $t_{0}$ represents the fixed pre-set time-step.
-Figure 5: The structure of the refiner module.
+where $t_{0}$ is the fixed pre-set time-step and $\mathcal{S}_{\mathrm{fixed}}$ denotes the vanilla SD time scheduler as shown in Eq. (9).
+Figure 4: The structure of the time scheduler module.
 It combines both the fixed scheduler and learnable scheduler to perform the denoising process.
@@
 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$
-DIV2K EDSR 24.64 22.13 0.6572 0.6056 0.5376 0.6219 0.3162 0.4184 34.15 23.28 0.3397 0.4193
-RCAN 24.34 22.06 0.6479 0.6034 0.5485 0.6307 0.3269 0.4349 31.57 23.18 0.3408 0.4091
-ESRGAN 23.15 21.15 0.5946 0.5808 0.4478 0.5788 0.2378 0.4237 59.80 43.40 0.6161 0.3725
+DIV2K ESRGAN 23.15 19.94 0.5946 0.5689 0.4478 0.5892 0.2378 0.4230 59.80 43.35 0.6161 0.3734
+StableSR 22.02 19.92 0.5983 0.5588 0.4980 0.5736 0.2078 0.2603 51.28 48.94 0.3780 0.3203
+S3Diff 20.22 17.81 0.5491 0.4601 0.4033 0.4895 0.1309 0.1810 64.37 67.92 0.6228 0.6991
 IRN 26.04 22.84 0.7020 0.6212 0.5153 0.6095 0.3041 0.4036 43.72 28.70 0.2799 0.2336
@@
 VQIR 23.91 22.02 0.6498 0.5823 0.3174 0.4568 0.1024 0.2663 64.04 58.21 0.6350 0.6293
-Ours 23.77 22.10 0.6675 0.5839 0.3026 0.4281 0.0824 0.1597 66.31 65.51 0.6443 0.6521
-Urban100 EDSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
-RCAN 19.96 18.14 0.5017 0.4452 0.5859 0.6880 0.3745 0.4808 40.31 24.46 0.3561 0.4130
-ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
+Ours 23.98 22.18 0.6736 0.5869 0.2979 0.4221 0.0886 0.1684 66.56 69.12 0.7189 0.7204
+CLIC2020 ESRGAN 25.43 22.47 0.7018 0.6804 0.4213 0.5429 0.2209 0.4011 57.80 41.28 0.6107 0.3558
+StableSR 24.11 22.24 0.7052 0.6726 0.4625 0.5288 0.1860 0.2282 52.69 49.80 0.3779 0.3325
+S3Diff 22.04 19.87 0.6615 0.5711 0.3968 0.4829 0.1336 0.1812 63.07 66.40 0.6001 0.6838
+IRN 28.55 24.93 0.7907 0.7228 0.4676 0.5528 0.2746 0.3776 47.66 27.81 0.2950 0.2495
+HCFlow 28.82 26.21 0.7949 0.7451 0.4542 0.5325 0.2635 0.3611 47.72 37.15 0.2733 0.2884
+VQIR 25.69 24.12 0.7363 0.6942 0.2854 0.4191 0.0901 0.2445 64.36 54.83 0.6105 0.6068
+Ours 26.04 24.26 0.7542 0.6903 0.2724 0.3911 0.0792 0.1513 65.35 69.10 0.6815 0.7271
+Urban100 ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
+StableSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
+S3Diff 17.60 15.74 0.4368 0.3419 0.4204 0.5282 0.1917 0.2573 71.01 69.56 0.6658 0.6776
 IRN 22.15 18.93 0.6155 0.4772 0.5114 0.6567 0.3338 0.4457 58.73 32.03 0.3217 0.2210
@@
 VQIR 20.27 18.44 0.5782 0.4427 0.3038 0.5045 0.1418 0.3290 71.56 61.64 0.6619 0.6568
-Ours 21.34 18.77 0.6464 0.4655 0.2601 0.4490 0.1107 0.2265 72.21 71.03 0.6773 0.6643
-CLIC2020 EDSR 26.33 23.92 0.7408 0.7062 0.5109 0.5742 0.3057 0.4024 33.17 21.77 0.3392 0.4307
-RCAN 26.37 24.04 0.7407 0.7072 0.5133 0.5777 0.3126 0.4151 31.59 22.07 0.3469 0.4180
-ESRGAN 25.43 23.48 0.7018 0.6991 0.4213 0.5347 0.2209 0.4014 57.80 40.98 0.6107 0.3537
-IRN 28.55 24.93 0.7907 0.7228 0.4676 0.5528 0.2746 0.3776 47.66 27.81 0.2950 0.2495
-HCFlow 28.82 26.21 0.7949 0.7451 0.4542 0.5325 0.2635 0.3611 47.72 37.15 0.2733 0.2884
-VQIR 25.69 24.12 0.7363 0.6942 0.2854 0.4191 0.0901 0.2445 64.36 54.83 0.6105 0.6068
-Ours 25.81 24.16 0.7497 0.6887 0.2777 0.3968 0.0754 0.1459 65.84 64.38 0.6198 0.6586
-DIV8K EDSR 26.56 24.48 0.6776 0.6472 0.5437 0.6048 0.3358 0.4325 30.83 22.30 0.3701 0.4686
-RCAN 26.63 24.57 0.6778 0.6481 0.5470 0.6085 0.3430 0.4449 29.93 22.52 0.3796 0.4502
-ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
+Ours 21.46 18.84 0.6495 0.4685 0.2577 0.4481 0.1154 0.2386 72.22 72.88 0.7155 0.6795
+DIV8K ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
+StableSR 25.90 23.55 0.6602 0.6258 0.4844 0.5497 0.2134 0.2623 47.15 46.97 0.4017 0.3557
+S3Diff 22.20 20.19 0.5903 0.5063 0.4163 0.4954 0.1524 0.1997 59.98 62.58 0.6351 0.6901
 IRN 28.95 25.46 0.7402 0.6636 0.4932 0.5870 0.3044 0.4066 42.61 27.76 0.3216 0.2741
@@
 VQIR 25.79 24.26 0.6889 0.6350 0.3221 0.4457 0.1066 0.2628 56.76 51.44 0.5769 0.6203
-Ours 25.93 24.15 0.7107 0.6302 0.3198 0.4376 0.0927 0.1798 57.75 58.89 0.6598 0.6461
-Figure 6: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results.
-The first row represents the original high-resolution images.
-Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
-Figure 7: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both the rescaled images and corresponding edge maps.
-Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
-Figure 8: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset
-Figure 9: Visual comparisons of downscaled LR images with sota methods.
-In this part, we fine-tune the denoising U-Net network using LoRA, while also jointly training the TPM and refiner module.
+Ours 26.16 24.37 0.7163 0.6337 0.3117 0.4358 0.0979 0.1888 58.54 63.34 0.6836 0.7229
+Finally, we fine-tune the denoising U-Net network and the VAE decoder using LoRA, while also jointly training the TPM and time scheduler module.
 To achieve a balance between fidelity and realism, we supervise the reconstruction results $\hat{x}$ using pixel loss and perceptual loss as follows:
-$\begin{split}\mathcal{L}_{\mathrm{enh.}}=&\underbrace{\left\|\hat{x}-x\right\|_{1}}_{\text{pixel loss}}+\\ &\underbrace{\lambda_{\mathrm{lpips}}\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\lambda_{\mathrm{dists}}\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)}_{\text{perceptual loss}}\end{split}$ (13)
+$\mathcal{L}_{\mathrm{enh.}}=$ $\left\|\hat{x}-x\right\|_{1}+$ (11)
+$\lambda_{\mathrm{pec.}}\left(\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)\right)$
 4 Experiments
@@
 4.1.1 Datasets and Evaluation Metrics.
-We train the proposed OSIRDiff using the DF2K dataset [1, 19].
+We train the proposed TADM using the DF2K dataset [1, 18].
 Considering that traditional benchmarks such as Set5, Set14, and BSD100 often have lower quality and resolution [5], they are not suitable for evaluating extreme rescaling methods.
-Therefore, we utilize four high-resolution datasets: DIV2K, Urban100 [7], CLIC2020 [32], and DIV8K [5] to evaluate models been compared.
+Therefore, we utilize four high-resolution datasets: DIV2K, CLIC2020 [25], Urban100 [6], and DIV8K [5] to evaluate models been compared.
 We adopt PSNR and SSIM [31] for quantitative evaluation.
@@
 4.1.2 Training Details
-OSIRDiff is built based on the SD 2.1-base model, and the training is divided into three stages.
-First, we train the DFRM using equation (8).
-Then we adopt equation (13) to jointly train the LoRA layers, the TPM and the refiner module.
-Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and refiner module using a combination of equations (8) and (13).
+TADM is built based on the SD 2.1-base model, and the training is divided into three stages.
+First, we train the DFRM using Eq.
+(8).
+Then we adopt Eq.
+(11) to jointly train the LoRA layers, the TPM and the time scheduler module.
+Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and scheduler module using a combination of Eq.
+(8) and (11).
 More training details are provided in the supplementary material.
 4.2 Results and Comparison with SOTA
-We compare the proposed OSIRDiff with three kinds of methods: (1) bicubic downscaling and SR, i.e., EDSR [19], RCAN [50] and ESRGAN [36]; (2) flow-based rescaling models, i.e., IRN [41] and HCFlow [18]; (3) prior-based extreme rescaling method VQIR [39].
+We compare the proposed TADM with three kinds of methods: (1) bicubic downscaling and SR, i.e., ESRGAN [29], StableSR [27] (based on multi-step diffusion) and S3Diff (based on single-step diffusion)[42]; (2) flow-based rescaling models, i.e., IRN [36] and HCFlow [17]; (3) prior-based extreme rescaling method VQIR [33].
 Some of them are proposed for 4$\times$ rescaling factors and we extended them to 16$\times$ or 32$\times$.
@@
 […]
+Figure 5: Qualitative comparisons of 16$\times$ (the 1st and 2nd rows) and 32$\times$ (the 3rd and 4th rows) rescaling results.
+Our method achieves higher semantic accuracy, such as more clearer text, more recognizable faces, and more realistic structures.
+Figure 6: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both the rescaled images and corresponding edge maps.
+Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
+Figure 7: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset. Our approach can still reconstruct correct semantic information even in such extreme scenarios.
 4.2.2 Qualitative Results
 In Fig.
-6, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors.
-Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as EDSR, RCAN, IRN, and HCFlow produce overly-smoothed results.
+5, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors.
+Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as IRN, and HCFlow produce overly-smoothed results.
 Although ESRGAN incorporates perceptual optimization it still cannot generate adequate details.
@@
 However, it struggles with faces and more complex structures.
-Thanks to the powerful prior provided by the pre-trained SD and the rich semantic information contained in the latent space of the pre-trained VAE, our method effectively addresses these issues.
-It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as faces, text and architectural structures.
+Although S3Diff can generate sharp edges and rich details, its reconstructed results have low fidelity and exhibit significant artifacts and color distortions.
+Thanks to the powerful prior provided by the pre-trained SD and the carefully designed architecture,, our method effectively addresses these issues.
+It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as text, faces, vehicles, and leaves.
 The quantitative experiments demonstrate that our model exhibits superior performance on the Urban100 [6] dataset, confirming that our method produces more accurate structural features.
 In this section, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig.
-7.
-Traditional regression models, such as EDSR [19], RCAN [50], IRN [41], and HCFlow [18], tend to generate overly-smooth results.
+6.
+Traditional regression models, such as IRN [36] and HCFlow [17], tend to generate overly-smooth results.
 GAN-based methods, including ESRGAN [29] and VQIR [33], generate more details, but their edge information lacks accuracy and regularity.
+For diffusion-based super-resolution methods, such as S3Diff [42], they are capable of generating sharp edges.
+However, due to the lack of information about downscaling, their results often exhibit lower fidelity, which manifests in distorted structures.
 In contrast, our approach, by leveraging the prior knowledge encapsulated in the SD model, is able to reconstruct more accurate and rich edge details.
 In Fig.
-8, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K [5].
-It is evident that our method is capable of recovering images with accurate semantics, even at extreme upscaling factors.
+7, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K.
+It is evident that our method is capable of recovering images with accurate semantics, even at extreme rescaling factors.
+For instance, our method is capable of accurately reconstructing electronic devices, doors, and chairs, while retaining rich details.
 To validate the effectiveness of our downscaling scheme, we conduct a visual comparison of the downscaled LR images with state-of-the-art (SOTA) methods, as shown in Fig.
-9.
+8.
 Overall, the LR images generated by our method achieve comparable or even superior visual quality to SOTA image rescaling methods.
@@
 In future work, we plan to further optimize the quality of the LR images.
-Figure 10: The validation loss curves when removing a specific component from the DFRM.
+Figure 8: Visual comparisons of downscaled LR images with sota methods.
 4.3 Ablation Study
-We conduct ablation experiments to validate the effectiveness of each component. All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K.
+We conduct ablation experiments to validate the effectiveness of each component.
+All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K for 16$\times$ rescaling task.
+Table 2: Ablation study of rescaling space and SD prior.
+Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
+$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
+$\checkmark$ $\times$ $\times$ 0.4675 0.3109
+$\checkmark$ $\times$ $\checkmark$ 0.2979 0.0886
 4.3.1 Effectiveness of Feature Rescaling
@@
 As shown in Fig.
-11, the rescaling operator in pixel space takes HR image as input and simultaneously outputs both the LR image and the rescaled image.
+9, the rescaling operator in pixel space takes HR image as input and simultaneously outputs both the LR image and the rescaled image.
 Then, to use SD for perceptual enhancement, we need to map the rescaled image to the latent space using a VAE encoder and perform the denoising process.
@@
 As shown in Fig.
-12, we conduct visual comparisons between rescaling in latent space and pixel space in two different scenarios.
+10, we conduct visual comparisons between rescaling in latent space and pixel space in two different scenarios.
 Our latent-space rescaling method is able to reconstruct more realistic textures and more accurate structural features.
@@
 Figure 10: Visual comparisons of our method with rescaling in pixel space and without the SD prior.
-Table 2: Ablation study of rescaling space and SD prior.
-Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
-$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
-$\checkmark$ $\times$ $\times$ 0.4675 0.3109
-$\checkmark$ $\times$ $\checkmark$ 0.3026 0.0824
 4.3.2 Necessity of SD Prior
@@
 Here, we compare the latent features before and after perceptual enhancement by decoding them to the pixel space using the VAE decoder, as shown in Fig.
-12.
+10.
 It can be seen that the latent features before perceptual optimization exhibit noticeable color shifts and blurriness when mapped to the pixel domain.
@@
 To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig.
-10.
+11.
 It can be observed that the introduction of the INN can reduce reconstruction error.
 Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as demonstrated by the lower guidance loss in Fig.
-10.
+11.
 As shown in Fig.
-13, the LR images exhibit significant noise and color distortions in the absence of pixel guidance.
+12, the LR images exhibit significant noise and color distortions in the absence of pixel guidance.
 However, the introduction of pixel guidance significantly alleviates these issues.
 Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
+Figure 11: The validation loss curves when removing a specific component from the DFRM.
 Figure 12: Visual comparisons of LR images with and without pixel guidance.
-Figure 14: Ablation study about time alignment.
 4.3.4 Effectiveness of Time Alignment
 To validate the effectiveness of our proposed time-step alignment strategy, we retrain our model using fixed time-steps.
-This involves removing the TPM and refiner modules and employing a manually selected time-step to perform a single denoising step based on equation (9).
+This involves removing the TPM and time scheduler modules and employing a manually selected time-step to perform a single denoising step based on Eq.
+(9).
 Following the setting of previous work [34], we select fixed time-steps 1 and 999 for comparison.
 As shown in Fig.
-14, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen.
+13, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen.
 In contrast, our time-step alignment mechanism can dynamically determine the level of generative capability based on the image content, thereby achieving optimal performance in both fidelity and perceptual quality.
-5 Discussions
-5.1 Comparison with JPEG Compression
+Figure 13: Ablation study about time alignment.
+Figure 14: Illustration of the tiled inference process and the predicted time-step map.
+In Fig.
+14, we present a schematic illustration of the tiled inference process along with the predicted time-step map.
+It can be observed that the time alignment mechanism effectively captures more complex regions in high-resolution images and assigns larger time steps to these areas.
+4.4 Discussions
+4.4.1 Details about Tiled Inference
+When TADM is employed for rescaling ultra-high-resolution images, it often necessitates dividing the input image into multiple patches for separate processing.
+In the main paper, this is referred to as the tiled inference strategy.
+In Algorithm 1, we elaborate on the process of tiled inference.
+Unlike works such as StableSR [27], our tiled inference algorithm can predict different time steps t for each image patch, thereby achieving dynamic allocation of generative capacity.
+Algorithm 1 Tiled Inference Process
+1:Input image $x$, latent encoder $\mathcal{E}$, decoupled feature rescaling module $\mathrm{DFRM}$, denoising U-Net $\epsilon_{\theta}$, time prediction module $\mathrm{TPM}$, time scheduler $\mathrm{TS}$, latent decoder $\mathcal{D}$, patch size $p$, stride length $s$
+2:Rescaled image $\hat{x}$, LR image $y$
+3:$z=\mathcal{E}(x)$triangleright latent encoding
+4:$\hat{z},y=\mathrm{DFRM}(z)$triangleright latent rescaling
+5:$[\hat{z}_{i}]=\mathrm{ToPatch}(\hat{z},p,s)$triangleright split $\hat{z}$ to patches
+6:$z_{0}\_{list}=[]$triangleright initialize list of $z_{0}$
+7:for $\hat{z}_{i}$ in $[\hat{z}_{i}]$ do
+8:     $t_{i}=\mathrm{TPM}(\hat{z}_{i})$triangleright time-step prediction
+9:     $\epsilon_{i}=\epsilon_{\theta}(\mathrm{patch},t_{i})$triangleright noise prediction
+10:     $z_{0}^{i}=\mathrm{TS}(\hat{z}_{i},\epsilon,t_{i})$triangleright denoising by time scheduler
+11:     $z_{0}\_{list}.\mathrm{append}(z_{0}^{i})$
+12:end for
+13:$z_{0}=\mathrm{Merge}(z_{0}\_{list})$triangleright merge patches of $z_{0}$
+14:$\hat{x}=\mathcal{D}(z_{0})$triangleright latent decoding
+15:return $\hat{x},y$triangleright output
+The aforementioned algorithm encompasses two hyperparameters: the patch size and the stride length, both defined in the latent space.
+If the patch size is set too large, although inference efficiency may improve, the prediction of time steps would become overly sparse, leading to performance degradation.
+Conversely, if the patch size is set too small, the computational load will be significantly higher, and there might be a misalignment between the inference size and the pre-training image size of Stable Diffusion.
+Therefore, to achieve a trade-off between performance and inference efficiency, we conduct experiments on the patch size, as shown in Fig.
+15.
+It can be observed that the model achieves optimal performance when the patch size is between 60 and 120.
+Consequently, we select 96 as the patch size in our work.
+For the stride length, we adopt the value from previous works [27] and directly set it to 64.
+In Fig.
+16, we present the time step predictions when employing different patch sizes during the tiled inference process.
+It can be observed that smaller patch sizes allow for finer-grained time step predictions.
+However, setting the size too small leads to misalignment with the pre-training image size of Stable Diffusion, resulting in increased computational load and performance degradation.
+In contrast, our selected patch size of 96 achieves a balance between performance and computational load, while also enabling relatively accurate prediction of the time step mask.
+Figure 15: Ablation study about patch size.
+Figure 16: The time step predictions at different patch sizes.
+4.4.2 Application to Image Compression
 Our model can be combined with existing lossless image compression methods to achieve extreme compression ratios.
-To validate this, we use the lossless compression image format PNG to compress LR images and compare its performance with JPEG compression at various bit-rates.
+To validate this, we use PNG to compress LR images and compare it with JPEG at various bit-rates.
 As shown in Fig.
-15, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression.
+17, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression.
 These comparisons demonstrate the potential of our model for image compression.
@@
 In Fig.
-16, we present the visual quality, the bpp and the LPIPS of the reconstructed images.
+18, we present the visual quality, the bpp and the LPIPS of the reconstructed images.
 Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics.
@@
 Figure 18: Visual comparisons between our model and JPEG compression. bpp/LPIPS of the reconstructed images are shown below the images.
-5.2 Time-step Distribution
-To overcome the non-uniform degradation introduced by the rescaling operation, we design a time-step alignment strategy to predict an optimal time-step $t$ based on the degraded latent variable $\hat{z}$, enabling dynamic allocation of generative capability.
-During the inference phase, our model can perform tiled inference, allowing it to predict different time-steps for various regions of the same image.
-We discuss the distribution of time-steps at both the single image and dataset scales.
-Figure 17: The time-steps $t$ predicted by the TPM on a single image, along with the corresponding latent variable rescaling error map $|z-\hat{z}|$.
-First, we visualize the time-steps corresponding to different regions of the image shown in Fig.
-17.
-We also present the latent variable rescaling error map $|z-\hat{z}|$ for comparison.
-It can be observed that in more complex regions, the rescaling operation introduces larger errors.
-Correspondingly, these areas also require larger time-steps.
-This is because complex regions necessitate greater generative capability for restoration.
-Figure 18: The distribution of time-steps predicted by the TPM on the DIV2K validation set at different rescaling scale factors. Smaller scale factor results in smaller time-steps.
-Next, we present the distribution of predicted time-steps at different rescaling scales on the DIV2K validation set shown in Fig.
-18.
-The vertical axis and the size of the bubbles represent the mean and variance of the time-steps, respectively.
-It can be observed that the mean of the predicted time-steps increases as the rescaling scale becomes larger.
-This is because larger rescaling scales result in greater image degradation, necessitating stronger generative capability for restoration.
-Additionally, we find that the predicted time-steps for the 16$\times$ rescaling exhibit the largest variance.
-This indicates that 16$\times$ rescaling is a task that balances both fidelity and generative capacity.
-In contrast, 8$\times$ rescaling tends to prioritize fidelity, while 32$\times$ rescaling leans more towards generative capabilities.
-6 Limitations
-Table 3: The performance comparison between our 4×4$\times$4 × rescaling model and VAE.
-Method PSNR $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
-VAE 29.00 0.1570 65.54 0.6234
-Ours-4×4$\times$4 × 29.32 0.1456 65.47 0.6313
-6.1 Upper Bound of OSIRDiff
-Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance ceiling is inherently constrained by the VAE itself.
-The VAE encoding is a lossy process and can be described as follows:
-$z=\mathcal{E}(x)$ (14)
-Here, $x\in R^{H\times W\times 3}$ represents original HR image and $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ represents corresponding latent features.
-We can measure the compression rate from the perspective of the number of bits:
-[…]
-$=\left(H\times W\times 3\right)\times 8$ (15)
-[…]
-$=\left(\frac{H}{8}\times\frac{W}{8}\times 4\right)\times 16$ (16)
-$\mathrm{Rate}$ $=\frac{\mathrm{Bits(z)}}{\mathrm{Bits(x)}}=\frac{1}{24}$ (17)
-It can be observed that VAE encoding can be regarded as a compression process with a compression ratio of 24.
-From the perspective of rescaling, this is equivalent to a rescaling operation with a scale factor of $S=\sqrt{24}$.
-Any rescaling scale smaller than $S$ will inevitably be limited by the reconstruction performance of VAE, as the VAE encoding process has already discarded significant amounts of information.
-Furthermore, we directly encode and decode input HR images using VAE without performing any rescaling operations, representing the upper performance bound of our model.
-Additionally, we train an OSIRDiff model for 4×4$\times$4 × rescaling and compare its performance with the reconstruction capability of the VAE, as illustrated in Table 3.
-It can be observed that the performance of our 4×4$\times$4 × rescaling model is very close to the reconstruction performance of the VAE, indicating that the rescaling performance at this scale is limited by the VAE reconstruction capacity.
-This also highlights a potential limitation of SD prior itself that it only demonstrates significant advantages when handling highly ill-posed tasks [8, 2, 46].
-Consequently, our proposed OSIRDiff is similarly affected by this limitation and performs well primarily at larger rescaling scales (>$S$).
-In the future, we could explore the incorporation of shallow features into the VAE decoding process to enhance its fidelity, thereby enabling the OSIRDiff framework to be more applicable to small-scale rescaling tasks.
-Table 4: Comparison of inference time and parameter sizes across methods.
-Time (s) Total Parameters (M) Trainable Parameters (M)
-IRN 0.359 34.2 34.2
-HCFlow 0.363 7.24 7.24
-VQIR 0.897 92.9 20.7
-Ours 0.711 1352.5 63.2
-6.2 Inference Time and Model Size
-In Table 4, we compare the inference speed and parameter sizes of different methods.
+4.4.3 Inference Efficiency Comparison
+In Table 3, we provide comparisons of inference time and memory consumption for representative methods.
 All tests are performed on an RTX 3090 using images with a resolution of $1024\times 1024$.
-Due to leveraging the SD architecture, our method has the largest total number of parameters and the highest number of trainable parameters.
-However, since our approach only requires a single denoising step, its inference time does not increase significantly compared to other methods.
-The inference speed of our method even surpasses that of the VQGAN-based [4] method, i.e.
-the VQIR.
-This is because, compared to VAE, the latent space of VQGAN has a higher number of channels.
-Additionally, the nearest-neighbor search process in VQGAN incurs significant computational overhead, adversely affecting its inference speed.
-In the future, we plan to explore architecture distillation techniques to further reduce the size of our image rescaling model and improve its inference speed.
+It can be observed that flow-based models (IRN and HCFlow) require less GPU memory consumption and enable fast inference.
+In contrast, StableSR, which is based on multi-step diffusion models, demands significantly more inference time and larger GPU memory consumption.
+Although VQIR requires less memory, the nearest neighbor search process in VQGAN results in longer inference time.
+For one-step diffusion-based method S3Diff, due to its degradation-guided LoRA which needs to dynamically predict LoRA parameters during inference, it requires longer ingerence time and higher memory consumption.
+In contrast, our proposed TADM demonstrates moderate efficiency and scalability compared to other methods.
+In the future, we can utilize distillation techniques to further reduce the model size.
+Table 3: Comparison of inference time and memory consumption.
+IRN HCFlow VQIR StableSR S3Diff Ours
+Inference Time (s) 0.359 0.363 0.897 135.2 2.186 0.711
+Memory Consumption (G) 0.498 0.098 0.344 6.380 5.025 2.598
+4.4.4 Upper Bound of TADM
+Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance upper boud is inherently constrained by the VAE.
+In Table 4, we present a performance comparison between VAE and two 4$\times$ rescaling models.
+For VAEs, we directly feed the high-resolution images to the encoder and decoder without performing any rescaling operations.
+This performance represents the upper bound of our method’s performance across all rescaling factors.
+It can be observed that VAE significantly lags behind the latter.
+In the future, we could explore integrating stronger VAEs, such as those from SDXL, SD3, thereby enabling the TADM to be more applicable to small-scale rescaling tasks.
+Table 4: Comparison of VAE and 4$\times$ rescaling methods.
+IRN HCFlow VAE
+PSNR 35.07 35.23 29.01
+SSIM 0.9318 0.9346 0.8046
 5 Conclusion
-We propose a novel framework called OSIRDiff to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling.
+We propose a novel framework called TADM to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling.
 Specifically, we design a DFRM to achieve bidirectional mapping between HR features and LR images.